### v1

# Provable Robustness in Multimodal Large Language Models via Feature Space Smoothing

Abstract
Multimodal large language models (MLLMs) exhibit strong capabilities across diverse applications, yet remain vulnerable to adversarial perturbations that distort their feature representations and induce erroneous predictions. To address this vulnerability, we propose the Feature-space Smoothing (FS) and theoretically prove that FS offers certified robustness on the feature representations of MLLMs. Specifically, FS transforms any feature encoder into a smoothed variant that is guaranteed to maintain a certified lower bound on the feature cosine similarity between clean and adversarial representations under $\ell_{2}$-bounded attacks. Moreover, we indicate that the value of this Feature Cosine Similarity Bound (FCSB) derived from FS can be improved by enlarging the defined Gaussian robustness score on the vanilla encoder. Building upon this, we introduce the Purifier and Smoothness Mapper (PSM), a plug-and-play module that improves the Gaussian robustness score of MLLMs and thus enhances their certified robustness under FS, without requiring any retraining on MLLMs. We demonstrate that the FS with PSM not only provides a strong theoretical robustness guarantee but also exhibits superior empirical performance compared to adversarial training. Extensive experiments across diverse MLLMs and downstream tasks indicate the effectiveness of the FS-PSM, reducing the Attack Success Rate (ASR) of various white-box attacks from nearly 90% to about 1%.
1 Introduction
The emergence of the Multimodal Large Language Models (MLLMs), such as GPT-5, Gemini 2.5, and Claude 3.7 Sonnet, has fundamentally reshaped existing working paradigms and significantly advanced societal productivity. Despite their remarkable capabilities across a broad spectrum of real-world tasks, these models still encounter critical safety challenges, such as adversarial vulnerabilities [51, 5, 23, 14].
Figure 1: Illustration of the FS-PSM, which guarantees that the cosine similarity of the adversarial and clean features extracted by MLLM’s encoder is larger than FCSB for robust predictions.
Adversaries can manipulate predictions of the MLLMs to a malicious state by injecting subtle and imperceptible perturbations to the clean inputs, exploiting the models’ insufficient local smoothness and uncontrolled Lipschitz continuity [10, 4, 12, 47].
Countermeasures towards those threats can be roughly classified into empirical defense and certified defense. Typical empirical approaches include adversarial training [27, 35, 41, 45, 2, 37, 28, 29] and input purification [27, 35, 41, 45, 2, 37, 28, 29]. Despite their demonstrated empirical effectiveness, the multimodal nature of MLLMs also poses a great challenge for existing adversarial defense methods. Since these models accept heterogeneous inputs across diverse domains, ensure a robust encoder via adversarial training that can generalize to various scenarios is challenging and computationally expensive. Moreover, these approaches lack formal robustness guarantees and remain susceptible to stronger or adaptive adversaries [38, 3, 44]. In contrast, certified approaches aim to guarantee that the model returns a constant prediction result within a certain range, usually a $\ell_{2}$ or $\ell_{\infty}$-norm constrained area [34, 43, 11, 15, 4, 47, 36]. However, most previous certified defense approaches [4, 47, 42, 11, 36] predominantly assume the prediction of one-dimensional outputs (e.g., class label), thereby limiting their applicability to more general tasks on MLLMs.
To address the aforementioned challenges, we propose the Feature-space Smoothing (FS), a provable defense method that offers certified robustness for the feature representations of MLLMs. FS smooths the vanilla feature encoder of the MLLM, and the resulted smoothed encoder is guaranteed to maintain a provable lower bound on the cosine similarity between clean and adversarial representations under $\ell_{2}$-norm bounded attacks.
Moreover, we demonstrate that the Feature Cosine Similarity Bound (FCSB) of the smoothed feature encoder is intrinsically determined by a Gaussian robustness score, which is defined to measure the prediction consistency of the vanilla feature extractor under Gaussian noise. Nonetheless, without Gaussian noise–augmented training, the Gaussian robustness of these MLLMs remains limited, leading to a correspondingly suboptimal FCSB. To address this and avoid the heavy computational cost of fine-tuning MLLMs, we propose a plug-and-play Purifier and Smoothness Mapper (PSM) that effectively enhance the Gaussian robustness score for the vanilla feature encoder in a training-free way. Specifically, the purifier operates prior to feature extraction and is trained to denoise Gaussian perturbations and maximize the Gaussian robustness score. Meanwhile, the smoothness mapper performs post-extraction refinement and is trained to preserve the feature distribution while further enhancing this score without fine-tuning the encoder. These two components of PSM work synergistically to enhance the certified robustness of MLLMs under FS.
To implement our PSM, we employ a pre-trained guided diffusion model as the Gaussian noise purifier and design a noise-aware residual network as the smoothness mapper. The PSM is optimized using the proposed utility–robustness loss and trained on data from diverse visual domains to enhance the Gaussian robustness while preserving the feature utility for the encoder. To comprehensively assess the performance, we evaluate our method against state-of-the-art (SOTA) adversarial attacks tailored for MLLMs under the white-box setting and compare it with strong adversarial training defense. Experimental results demonstrate that our FS with PSM enhanced provides strong protection for various MLLMs under diverse downstream tasks. Overall, the main contributions of this work are:
  • We propose the Feature-space Smoothing (FS) to turn any feature encoder of MLLMs into a smoothed version, and theoretically prove that the smoothed encoder maintains a certified lower bound on the feature cosine similarity between clean and adversarial representations.
  • We propose a plug-and-play Purifier and Smoothness Mapper (PSM) that effectively enhances the value of the FCSB of the smoothed encoder in a training-free manner.
  • We conduct extensive experiments demonstrating that integrating our proposed FS-PSM greatly enhances the adversarial performance of various MLLMs and dramatically reduces the ASR under various white-box attacks.
2 Related Work
Adversarial attacks on MLLMs. While MLLMs continue to achieve remarkable performance across diverse applications, extensive works [31, 5, 51, 14, 23, 39, 49, 50, 48] have exposed their adversarial vulnerabilities, raising serious safety concerns. Early work, such as AttackVLM [51], explores transferable attacks by disrupting the feature representations of CLIP [33] and BLIP [19], showing strong adversarial transferability among open-source models but limited effectiveness against closed-source commercial systems. More recent approaches, such as M-Attack [23] and FOA-Attack [14], further advance this direction by leveraging multi-extractor ensembles and feature-space alignment, achieving over 90% targeted attack success rates on image-captioning tasks against powerful closed-sourced commercial MLLMs (e.g., ChatGPT-4o). This highlights an urgent need for trustworthy defenses that provide effective and provable protection for MLLMs.
Adversarial defense on MLLMs. Adversarial defense methods can be broadly classified into empirical and provable approaches. Empirical defenses for MLLMs mainly include adversarial training [27, 35, 41, 45, 2, 37, 28, 29], which enhances robustness by augmenting training data with adversarial examples, and input purification [30, 17, 22], which employs generative mechanisms such as diffusion models or autoencoders to recover clean inputs prior to inference. Recent studies [28, 29, 37] have revealed that utilization of adversarially trained CLIP-feature encoders can enhance adversarial robustness for MLLMs. However, these methods lack formal robustness guarantees and remain vulnerable to adaptive and unseen threats [38, 3, 44]. Moreover, adversarial training demands costly retraining and often leads to degradation in clean performance.
Alternatively, certified defenses aim to provide mathematically provable robustness guarantees. The use of Gaussian smoothing for certified robustness was initially introduced for classification models [4, 18, 16], yet its theoretical formulation is restricted to one-dimensional outputs, limiting its applicability to tasks such as auto-regression or multimodal generation. To overcome these limitations and ensure trustworthy protection for MLLMs, we propose the Feature-space Smoothing (FS), a defense mechanism that establishes provable adversarial robustness for MLLMs.
3 Feature Space Smoothing
3.1 Preliminary
Let $\mathcal{F}$ denote a general deep learning model consisting of a feature extractor $f_{e}:{\bm{x}}\rightarrow{\bm{z}}$ that maps the input $\bm{x}$} to a feature representation $\bm{z}$}, and a predictor $f_{d}:{\bm{z}}\rightarrow{\bm{y}}$ that produces the final output $\bm{y}$}. Let $\mathcal{L}$ denote the general loss function (e.g., cross-entropy) that measures the discrepancy between the model’s output and the ground truth.
Adversarial attacks: Let $\mathcal{B}_{\epsilon}({\bm{x}})=\left\{{{\bm{x}}^{\prime}:{{\left\|{{\bm{x}}^{\prime}-{\bm{x}}}\right\|}_{p}}\leq\epsilon}\right\}$ be an $\ell_{p}$-norm ball centered at the input $\bm{x}$}, where $\epsilon$ is a pre-defined perturbation bound. For each input $\bm{x}$}, the adversarial attacks aim to find an adversarial input ${\bm{x}}^{\prime}={\bm{x}}+{\bm{\delta}}$ that misleads the model by solving:
$\vskip-2.84526pt\mathop{\max}\limits_{{{\bm{x}}}+{\bm{\delta}}\in\mathcal{B}_{\epsilon}({\bm{x}})}\mathcal{L}\left({\mathcal{F}\left({{{\bm{x}}}}\right),\mathcal{F}\left({{{\bm{x}}}+{\bm{\delta}}}\right)}\right).\vskip-2.84526pt$ (1)
Adversarial effects on feature representations: While adversarial attacks primarily aim to alter the model’s predictions, numerous studies [14, 46, 20, 13, 21, 23, 40, 32, 8] have shown that successful attacks typically induce substantial distortions in the model’s feature representations. Let $\bm{x}$$^{\prime}$ denote the adversarial example with adversarial feature $\bm{z}$$^{\prime}$. Let ${\bm{z}}^{t}$ be the adversarial targeted feature with malicious semantic meaning. The attack generally leads to $\max\mathcal{L}\left({{\bm{z}}^{\prime},{\bm{z}}}\right)$ for untargeted attacks and $\min\mathcal{L}\left({{\bm{z}}^{\prime},{\bm{z}}^{t}}\right)$ for targeted attacks. Thus, ensuring a robust feature encoder that $\min\mathcal{L}\left({{\bm{z}}^{\prime},{\bm{z}}}\right)$ is crucial for the trustworthy prediction.
Randomized smoothing: Consider a $k$ classes classification problem with the input ${\bm{x}}\in{\mathbb{R}^{d}}$ and the label ${\bm{y}}\in\mathcal{Y}=\left\{{{c_{1}},\ldots,{c_{k}}}\right\}$. Randomized Smoothing (RS) first corrupts each input $\bm{x}$ by adding the Gaussian noise $\bm{\varepsilon}\sim\mathcal{N}(0,{\sigma^{2}}\bm{I})$. Then it turns an arbitrary base classifier $\mathcal{F}$ into a smoothed version $\hat{\mathcal{F}}$ that possesses $\ell_{2}$} certified robustness guarantees. The smoothed classifier $\hat{\mathcal{F}}$ returns whichever the class the base classifier $\mathcal{F}$ is most likely to return among the distribution ${\bm{x}}+\bm{\varepsilon}\sim\mathcal{N}({\bm{x}},{\sigma^{2}}\bm{I})$, which is:
$\begin{array}[]{l}\hat{\mathcal{F}}({\bm{x}})=\mathop{\arg\max}\limits_{c\in\mathcal{Y}}{\mathbb{P}}(\mathcal{F}({\bm{x}}+\bm{\varepsilon})=c).\vskip-5.69054pt\end{array}$ (2)
RS then guarantees a certified radius $\mathcal{R}$ for this smoothed classifier $\hat{\mathcal{F}}$. For any perturbation $\delta$ satisfying $\left\|\delta\right\|_{2}\leq\mathcal{R}$, the smoothed classifier is guaranteed to return a robust prediction that makes $\hat{\mathcal{F}}({\bm{x}}+\delta)=F(x)$.
Limitations for RS. While previous RS provides effective certified protection for classification models, it suffers from two limitations. First, the theoretical framework of RS inherently restricts its certification to classification tasks (e.g. possibility of input $\bm{x}$} belongs to a certain class $c$). Second, estimating $\mathbb{P}(\mathcal{F}({\bm{x}}+\bm{\varepsilon})=c)$ in Equation 2, incurs substantial computational overhead, since each estimation requires multiple forward passes through the entire model.
3.2 Certified Bound via Feature Space Smoothing
Considering the limitations inherent in the RS, we introduce the Feature-space Smoothing (FS). By turning any feature encoder $f_{e}$ into a smoothed version $\hat{f_{e}}$, FS theoretically guarantees that $\hat{f_{e}}$ maintains a certified lower bound on the cosine similarity between clean and adversarial representations under $\ell_{2}$-norm constrained perturbations.
Smoothed feature encoder. For any feature encoder $f_{e}:{\bm{x}}\rightarrow{\bm{z}}$, where $\bm{z}$} is the representation normalized into the $l_{2}$ unit sphere, FS defines the smoothed encoder $\hat{f}_{e}(\bm{x})$ as:
$\hat{f}_{e}(\bm{x})=\mathbb{E}_{\bm{\varepsilon}\sim\mathcal{N}(0,I)}[f_{e}(\bm{x}+\bm{\varepsilon})]$
$=\frac{1}{(2\pi)^{d/2}}\int_{\mathbb{R}^{d}}f_{e}(\bm{x}+\bm{\varepsilon})\exp\!\left(-\tfrac{1}{2}\|\bm{\varepsilon}\|^{2}\right)d\bm{\varepsilon},\vskip-2.84526pt$ (3)
where $I$ denotes the $d\times d$ identity and $\|\cdot\|$ is the Euclidean norm. $\bm{\varepsilon}\sim\mathcal{N}(0,I)$ denotes Gaussian noise with zero mean and standard deviation one. The smoothed feature encoder outputs the expectation of feature representations over the Gaussian distribution $\mathcal{N}({\bm{x}},I)$.
Figure 2: The training framework of the PSM. The purifier performs pre-processing and the smoothness mapper refines post-extracted features to enhance the Gaussian robustness. Parameters of MLLMs are frozen, and the purifier and mapper are optimized with $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\mathcal{M}}$. For evaluation, the input $\bm{x}$} is replaced with adversarial input $\bm{x}$$^{\prime}$ and the forward calculation marked by orange color will be removed.
Gaussian robustness score. Define ${S_{\bm{x_{t}}}}({\bm{x}})$ as the score function that evaluates feature discrepancy between an input $\bm{x}$} and a targeted example ${\bm{x}}_{t}$, which is:
$\vskip-5.69054ptS_{{\bm{x}}_{t}}({\bm{x}})=\tfrac{1}{2}\Big(1+\mathrm{Cos}\big(f_{e}({\bm{x}}),\,f_{e}({\bm{x}}_{t})\big)\Big)$ (4)
where ${Cos}(\cdot,\cdot)$ denotes the cosine similarity and $S_{{\bm{x}}_{t}}({\bm{x}})\in[0,1]$. The Gaussian robustness score $\hat{S}({\bm{x}})$ is defined as:
$\hat{S}({\bm{x}})=\mathop{\mathbb{E}}\limits_{\bm{\varepsilon}\sim\mathcal{N}(0,I)}\big[S_{{\bm{x}}}({\bm{x}}+\bm{\varepsilon})\big]$ (5)
$=\tfrac{1}{2}\!\left(1+\mathop{\mathbb{E}}\limits_{\bm{\varepsilon}\sim\mathcal{N}(0,I)}\!\left[\mathrm{Cos}\big(f_{e}({\bm{x}}+\bm{\varepsilon}),\,f_{e}({\bm{x}})\big)\right]\right),$
The score $\hat{S}({\bm{x}})$ evaluates the expected cosine similarity between the feature representation of a Gaussian-perturbed input ${\bm{x}}+\bm{\varepsilon}$ and that of the clean input $\bm{x}$}, characterizing the Gaussian robustness of the vanilla feature encoder ${f}_{e}({\bm{x}})$. Notably, we prove that this score $\hat{S}({\bm{x}})$ preserves a good Lipschitz property, which serves as a theoretical foundation for proving the certified robustness for $\hat{f}_{e}({\bm{x}})$.
Lemma 1 (Lipschitz property for the Gaussian robustness score).
Let $\Phi(a)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{a}\exp(-\frac{1}{2}s^{2})ds$ be a standard Gaussian cumulative distribution function and $\Phi^{-1}$ be its inverse. For any feature encoder $f_{e}:{\bm{x}}\rightarrow{\bm{z}}$, the mapping ${\bm{x}}\rightarrow\Phi^{-1}(\hat{S}(\bm{x}))$ is 1Lipschitz1-Lipschitz.
The proof of Lemma 1 is in the supplementary material, Section S.1. It implies that the mapping from $\bm{x}$} to $\Phi^{-1}(\hat{S}({\bm{x}}))$ exhibits strong adversarial robustness, as it satisfies a $1$-Lipschitz constraint. Generally, $\hat{S}({\bm{x}})$ measures the Gaussian robustness of the vanilla feature encoder $f_{e}(x)$. We then prove that this score $\hat{S}({\bm{x}})$ fundamentally determines the value of the certified robustness bound of its smoothed encoder $\hat{f}_{e}(x)$. The following Theorem establishes an explicit relationship between this score and the lower bound on the adversarial feature cosine similarity.
Theorem 1 (Certified lower bound on the adversarial feature cosine similarity).
For any feature encoder ${f}_{e}$ and its smoothed version $\hat{f}_{e}$, let $\bm{x}$} and $\bm{x}$$^{\prime}$ be clean and adversarial inputs with $\|{\bm{x}}^{\prime}-{\bm{x}}\|\leq\epsilon$. The cosine similarity between the adversarial feature $\hat{f}_{e}({\bm{x}}^{\prime})$ and clean feature ${f}_{e}({\bm{x}})$ satisfies: $\mathrm{Cos}(\hat{f}_{e}({\bm{x}}^{\prime}),f_{e}({\bm{x}}))\geq 2{\Phi}\left({\Phi^{-1}}(\hat{S}({\bm{x}}))-\epsilon\right)-1$.
Denote $2{\Phi}\left({\Phi^{-1}}(\hat{S}({\bm{x}}))-\epsilon\right)-1$ as the Feature Cosine Similarity Bound (FCSB). Theorem 1 reveals that:
  • By FS, we can turn any given feature encoder $f_{e}$ into a smoothed version $\hat{f}_{e}$ that maintains a FCSB between the adversarial and clean feature representations.
  • By maximizing the robustness score $\hat{S}({\bm{x}})$ of the given feature encoder $f_{e}$, we can effectively enhance the value of FCSB derived on its smoothed version $\hat{f}_{e}$.
The proof of Theorem 1 is in the supplementary material, Section S.1.
Corollary 1 (Certified radius $\mathcal{R}$ for adversarial cosine similarity $\geq$ 0.5).
Let $\bm{x}$} be the clean input, and $\bm{x}$$^{\prime}$ be the adversarial input. Then $\mathrm{Cos}(\hat{f}_{e}(x^{\prime}),{f}_{e}(x))\geq 0.5$, for all $\bm{x}$$^{\prime}$ with $\|{\bm{x}}^{\prime}-{\bm{x}}\|_{2}\leq\mathcal{R}$, where:
$\vskip-2.84526pt\mathcal{R}=\Phi^{-1}\!\big(\hat{S}({\bm{x}})\big)-\Phi^{-1}(0.75).$ (6)
Building upon Theorem 1, Corollary 1 establishes a certified radius $\mathcal{R}$ for the smoothed feature encoder. For any adversarial perturbation satisfying $\|\delta\|_{2}\leq\mathcal{R}$, the FCSB of the smoothed encoder is guaranteed to remain above 0.5.
Why feature-space smoothing for MLLMs? Unlike previous smoothing methods, performing smoothing in the feature space offers several advantages:
  • Efficiency: Compared to the previous RS that smooths the entire model $\mathcal{F}$, smoothing the feature encoder $f_{e}$, which is substantially lighter than the whole MLLM $\mathcal{F}$, significantly reduces the computational time of multiple forward calculations required by smoothing.
  • Generality: The FS provides certified robustness at the feature representation level, making it applicable to various downstream tasks (e.g. image captioning, classification and visual question answering).
  • Effectiveness: Since feature representations play a critical role in the final prediction, guaranteeing a trustworthy feature representation effectively improves the prediction reliability and robustness under adversarial attacks.
4 Purifier and Smoothness Mapper
Theorem 1 and Corollary 1 reveal an intriguing robustness property of the smoothed feature encoder $\hat{f}_{e}$. However, MLLMs’ feature encoders $f_{e}$ generally exhibits limited Gaussian robustness, which restricts the value of FCSB derived on $\hat{f}_{e}$. One solution is to estimate the smoothness score $\hat{S}({\bm{x}})$ through Monte Carlo sampling and training $f_{e}$ to maximize this score via gradient backpropagation. However, this could reduce the adaptability and practicality of our FS protection, as fine-tuning and re-aligning the LLM with multi-modal feature encoders is highly complex and costly. This thereby motivates us to propose the Purifier and Smoothness Mapper (PSM), a plug-and-play module that can be seamlessly integrated with MLLMs to enhance its Gaussian robustness score $\hat{S}(x)$.
The training framework of our proposed PSM is shown in Figure 2. The purifier $\mathcal{P}$ operates before feature extraction to denoise the Gaussian perturbations, and the smoothness mapper $\mathcal{M}$ performs post-extraction to do the feature refinement. Those two modules work together to enhance the Gaussian robustness of the given feature encoder.
4.1 Noise purifier
To denoise the Gaussian noise, $\mathcal{P}$ is trained to minimize the reconstruction loss $l_{mse}$, defined as:
$\vskip-2.84526ptl_{mse}=\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}{\bm{x}}\sim\mathcal{D},~\bm{\varepsilon}\sim\mathcal{N}(0,\sigma^{2}I)\end{subarray}}\|{\bm{x}}-\mathcal{P}({\bm{x}}+\mathbf{\varepsilon})\|,$ (7)
where $\mathcal{D}$ represents the data distribution. Meanwhile, to further enhance robustness score after plugging $\mathcal{P}$, we also introduce a robustness loss $l^{\mathcal{P}}_{rb}$, defined as:
$\vskip-2.84526ptl^{\mathcal{P}}_{\mathrm{rb}}=\mathop{\mathbb{E}}\limits_{{\bm{x}}\sim\mathcal{D},~\bm{\varepsilon}\sim\mathcal{N}(0,\sigma^{2}I)}\big[\mathrm{Cos}\big(f_{e}(\mathcal{P}({\bm{x}}+\bm{\varepsilon})),f_{e}({\bm{x}})\big)\big],$ (8)
which encourages feature consistency between the purified and clean representations. Practically, we adopt the ImageNet-pretrained guided-diffusion [7] as our purifier $\mathcal{P}$ and perform one-step diffusion for denoising in our main experiments. We then fine-tune this model on the dataset $\mathcal{D}$ using the purifier loss $\mathcal{L}_{\mathcal{P}}$, which is:
$\vskip-2.84526pt\mathcal{L}_{\mathcal{P}}=l_{\mathrm{diff}}+\lambda_{1}l^{\mathcal{P}}_{\mathrm{rb}}+\lambda_{2}l_{mse},$ (9)
where $l_{\mathrm{diff}}$ denotes the original diffusion loss, and $\lambda_{1}$ and $\lambda_{2}$ are the weighting coefficients. More details of the fine-tuning process on $\mathcal{P}$ can be found in the supplementary material, Section S.2.
4.2 Residual smoothness mapper
For the residual smoothness mapper $\mathcal{M}$, we utilize a noise-aware residual module to enhance feature robustness while preserving its statistical distribution. The main process of this mapper can be formulated as:
$\vskip-8.53581pt\tilde{{\bm{z}}}_{m}=\tilde{{\bm{z}}}+\mathcal{M}(\tilde{{\bm{z}}},\sigma)=\tilde{{\bm{z}}}+\sum_{i=0}^{k-1}m_{i}\left(\tilde{{\bm{z}}}_{i},\sigma\right),$ (10)
where $\tilde{{\bm{z}}}=f_{e}(\mathcal{P}({\bm{x}}+\mathbf{\varepsilon}))$ denotes the purified feature representations and $\tilde{{\bm{z}}}_{i+1}=m_{i}(\tilde{{\bm{z}}}_{i},\sigma)$ is the intermediate output with $\tilde{{\bm{z}}}_{0}=\tilde{{\bm{z}}}$. $\sigma$ is the noise strength that adaptively controls the output magnitude of the mapper. $k$ is the number of blocks ($k=3$ in our experiments unless otherwise specified), and each block $m_{i}(\cdot)$ contains multi-head attention, depthwise convolution, and MLP branches to refine the purified representation. To enhance the Gaussian robustness of the refined representation, we introduce the mapper robustness loss $l^{\mathcal{M}}_{\mathrm{rb}}$, defined as:
$\vskip-2.84526ptl^{\mathcal{M}}_{\mathrm{rb}}=\mathop{\mathbb{E}}\limits_{{\bm{x}}\sim\mathcal{D},~\bm{\varepsilon}\sim\mathcal{N}(0,\sigma^{2}I)}\big[\mathrm{Cos}\big(\tilde{{\bm{z}}}_{m},f_{e}({\bm{x}})\big)\big],\vskip-2.84526pt$ (11)
which encourages feature consistency between the refined and clean representations. Meanwhile, to ensure that the refined feature preserves the statistical characteristics of the clean feature, we introduce two regularization terms: the identical loss $l_{\mathrm{id}}$ and the statistical loss $l_{\mathrm{stats}}$, defined as:
$\vskip-2.84526pt\left\{\small\begin{aligned} &l_{\mathrm{stats}}=\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}{\bm{x}}\sim\mathcal{D},\\ \bm{\varepsilon}\sim\mathcal{N}(0,\sigma^{2}I)\end{subarray}}\frac{1}{D}\sum_{d=1}^{D}\Big[(\mu_{\tilde{{\bm{z}}}_{m}}^{(d)}-\mu_{{\bm{z}}}^{(d)})^{2}+(\sigma_{\tilde{{\bm{z}}}_{m}}^{(d)}-\sigma_{{\bm{z}}}^{(d)})^{2}\Big],\\ &l_{\mathrm{id}}=\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}{\bm{x}}\sim\mathcal{D}\end{subarray}}\|\mathcal{M}(\tilde{{\bm{z}}},0)\|_{2}^{2}.\end{aligned}\right$. (12)
where $\tilde{{\bm{z}}}_{m},{\bm{z}}\in\mathbb{R}^{B\times L\times D}$ with bath size $B$, token number $L$ and feature dimension $D$. The $l_{\mathrm{stats}}$ enforces consistency between the element-wise mean $\mu^{d}$ and standard deviation $\sigma^{d}$ of two representations, thereby preserving the statistical characteristics. Meanwhile, the identity loss $l_{\mathrm{id}}$ constrains the mapping network $\mathcal{M}$ when the noise strength $\sigma=0$, promoting stability and preventing undesired distortions on clean inputs. The overall training loss $\mathcal{L}_{\mathcal{M}}$ is defined as:
$\vskip-2.84526pt\mathcal{L}_{\mathcal{M}}=l^{\mathcal{M}}_{\mathrm{rb}}+\lambda_{3}l_{\mathrm{stats}}+\lambda_{4}l_{\mathrm{id}},$ (13)
where $\lambda_{3}$ and $\lambda_{4}$ are the weighting coefficients. More details of the training process on the residual smoothness mapper can be found in the supplementary material, Section S.2.
4.3 Further discussion on PSM
Certified robustness for the encoder $f_{e}$ with PSM. Let $f^{\prime}_{e}$ denote the feature encoder integrated with the proposed PSM. Then the forward process can be formulated as:
$\vskip-2.84526pt{f}^{\prime}_{e}(\bm{x}+\bm{\varepsilon})=f_{e}(\mathcal{P}({\bm{x}}+\bm{\varepsilon}))+\mathcal{M}(\tilde{{\bm{z}}},\sigma).$ (14)
Under this condition, the smoothed feature encoder and smoothness score are defined as:
$\hat{f}^{\prime}_{e}(\bm{x})=\frac{1}{(2\pi)^{d/2}}\int_{\mathbb{R}^{d}}f^{\prime}_{e}(\bm{x}+\bm{\varepsilon})\exp\left(-\tfrac{1}{2}\|\bm{\varepsilon}\|^{2}\right)d\bm{\varepsilon},$ (15)
$\hat{S}^{\prime}(\bm{x})=\tfrac{1}{2}\left(1+\mathop{\mathbb{E}}\limits_{\bm{\varepsilon}\sim\mathcal{N}(0,\sigma^{2}I)}\left[{Cos}\big(f^{\prime}_{e}({\bm{x}}+\bm{\varepsilon}),\,f_{e}({\bm{x}})\big)\right]\right),$
Where $\hat{S}^{\prime}(\bm{x})\in[0,1]$. We can prove that the Lipschitz property derived in Lemma 1 still holds, and the theoretical bound in Section 3 remains valid. Then, utilizing Theorem 1 and Corollary 1, we can derive the certified lower bound on $\mathrm{Cos}(\hat{f}^{\prime}_{e}(\bm{x^{\prime}}),f_{e}(\bm{x}))$ for any adversarial input $\bm{x}$$^{\prime}$.
Algorithm 1 Training algorithm of PSM
1:Input: dataset $\mathcal{D}$, feature encoder $f_{e}$, purifier $\mathcal{P}$, mapper $\mathcal{M}$, sampling number $n_{0}$, noise standard deviation $\sigma$, and loss weights $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$.
2:Output: trained purifier $\mathcal{P}$ and mapper $\mathcal{M}$.
3:// Stage 1: Fine-tune the purifier $\mathcal{P}$
4:for each batch ${\bm{x}}\sim\mathcal{D}$ do
5:  ${\bm{z}}=f_{e}({\bm{x}})$
6:  for $i=1,\dots,n_{0}$ do
7:    Sample $\bm{\varepsilon}^{i}\sim\mathcal{N}(0,\sigma^{2}I)$
8:    $\tilde{{\bm{x}}}^{i}=\mathcal{P}({\bm{x}}+\bm{\varepsilon}^{i})$, and $\tilde{{\bm{z}}}^{i}=f_{e}(\tilde{{\bm{x}}}^{i})$
9:  Compute $l^{\mathcal{P}}_{\mathrm{rb}}$ and $l_{\mathrm{mse}}$ using Equations 78.
10:  Update $\mathcal{P}$ by gradient descent on $\nabla_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}$
11:// Stage 2: Train the smoothness mapper $\mathcal{M}$
12:for each batch ${\bm{x}}\sim\mathcal{D}$ do
13:  ${\bm{z}}=f_{e}({\bm{x}})$
14:  for $i=1,\dots,n_{0}$ do
15:    Sample $\bm{\varepsilon}^{i}\sim\mathcal{N}(0,\sigma^{2}I)$
16:    $\tilde{{\bm{x}}}^{i}=\mathcal{P}({\bm{x}}+\bm{\varepsilon}^{i})$, and $\tilde{{\bm{z}}}^{i}_{m}=f_{e}(\tilde{{\bm{x}}}^{i})+\mathcal{M}(\tilde{{\bm{z}}}^{i},\sigma)$
17:  Compute $l^{\mathcal{M}}_{\mathrm{rb}}$, $l_{\mathrm{id}}$ and $l_{\mathrm{stats}}$ using Equations 1112.
18:  Update $\mathcal{M}$ by gradient descent on $\nabla_{\mathcal{M}}\mathcal{L}_{\mathcal{M}}$
Table 1: Experimental results on adversarial robustness of different defense methods and MLLMs on image captioning tasks. Values in parentheses denote the average textual similarity measured by GPTScore. The overall best results are shown in bold, and the best results without smoothing are underlined, highlighting the significant performance gain introduced by our proposed FS-PSM.
Model Method M-Attack [23] FOA [14] AttackVLM [51]
[…]
LLaVA- 1.5-7B Org. 0.385 1% (0.06) 93% (0.54) 0.388 1% (0.049) 94% (0.578) 0.430 3% (0.089) 88% (0.490)
Smoothed org. 0.650 82% (0.694) 2% (0.028) 0.652 87% (0.735) 1% (0.021) 0.660 89%(0.727) 2% (0.018)
FARE [37] 0.588 44% (0.409) 24% (0.097) 0.504 19% (0.197) 51% (0.222) 0.499 18%(0.190) 44%(0.221)
Smoothed FARE 0.716 59% (0.501) 14% (0.069) 0.656 35% (0.327) 29% (0.153) 0.729 60% (0.503) 11% (0.061)
TeCoA [29] 0.720 51% (0.458) 16% (0.081) 0.674 21% (0.236) 37% (0.187) 0.606 17% (0.179) 43% (0.218)
Smoothed TeCoA 0.805 59% (0.498) 14% (0.056) 0.789 39% (0.364) 20% (0.094) 0.755 42%(0.358) 18% (0.104)
Open Flamingo -9B Org. 0.351 1%(0.089) 86%(0.583) 0.347 1%(0.091) 87%(0.569) 0.442 16%0.222) 59%(0.379)
Smoothed org. 0.673 73%(0.655) 1%(0.063) 0.672 69%(0.623) 0%(0.061) 0.703 73%(0.687) 0%(0.055)
FARE [37] 0.588 36%(0.407) 17%(0.145) 0.504 20%(0.249) 40%(0.240) 0.499 15%(0.243) 40%(0.247)
Smoothed FARE 0.716 55%(0.529) 9%(0.088) 0.656 40%(0.414) 17%(0.125) 0.729 50%(0.488) 4%(0.089)
TeCoA [29] 0.720 51%(0.472) 13%(0.116) 0.674 21%(0.240) 30%(0.210) 0.606 19%(0.244) 34%(0.225)
Smoothed TeCoA 0.805 56%(0.518) 5%(0.090) 0.789 50%(0.454) 8%(0.095) 0.755 38%(0.398) 12%(0.126)
Training algorithm and dataset construction. The training procedure of PSM is summarized in Algorithm 1, where the purifier $\mathcal{P}$ and the smoothness mapper $\mathcal{M}$ are trained sequentially via a two-stage manner. The expectation over Gaussian perturbations is approximated by Monte Carlo sampling with $n_{0}$ samples drawn from $\mathcal{N}(0,\sigma^{2}I)$, where we set $n_{0}=8$ in our training to balance efficiency and estimation accuracy. Notably, all modules can be trained in a self-supervised fashion, without relying on task-specific labels. Thus, to promote generalization across diverse visual domains and ensure compatibility with various MLLMs’ downstream tasks, we collect the training dataset $\mathcal{D}$ comprising 5,000 images from heterogeneous domains, including medical, cartoon, and natural images.
Table 2: The average FCSB at different predetermined adversarial bounds $\epsilon$ and the average certified radius $\mathcal{R}$ for $\mathrm{FCSB}\geq 0.5$.
Encoder $\sigma$ Avg. FCSB at different adv-bound $\epsilon$ Avg. $\mathcal{R}$
0.125 0.25 0.375 0.50 0.75
CLIP-B16 0.25 0.828 0.623 0.313 -0.06 / 0.31
[…]
CLIP-B16 0.50 0.717 0.586 0.439 0.253 -0.132 0.33
[…]
Certified robustness bound of FS-PSM. To evaluate the effectiveness of the proposed PSM module, we adopt CLIP-B16 [33] as the vanilla feature encoder $f_{e}$ and assess the certified robustness bound of both the original smoothed encoder $\hat{f}_{e}$ and its PSM-enhanced counterpart $\hat{f}^{\prime}_{e}$ under the FS. We report both the certified FCSB under different adversarial constraints $\epsilon$ and the certified radius $\mathcal{R}$ for $\mathrm{FCSB}\geq 0.5$. The results are shown in Table 2. For implementation, we randomly sample 100 images from the ImageNet dataset [6] and approximate the Gaussian expectation using n=1,000n=1{,}000 Monte Carlo samples. As shown in Table 2, integrating PSM consistently enhances both the certified FCSB across a range of perturbation magnitudes and the average certified radius $\mathcal{R}$, indicating substantial improvements in certified robustness. To further assess the practical protection of our FS-PSM, we then conduct extensive experiments across multiple MLLMs and downstream tasks in the next section, evaluating their performance before and after smoothing under various strong white-box attacks.
5 Experiments
5.1 Experimental setup
Evaluated models and tasks. Since the proposed FS requires access to the forward feature computation process of MLLMs, we primarily validate its effectiveness on open-sourced MLLMs, including LLaVA-V1.5-7B [24, 25] and OpenFlamingo9B [1]. We comprehensively assess the performance of plugging the FS-PSM under adversarial conditions across multiple downstream tasks, including:
  • Image captioning: Following the setup in [23, 14], we randomly take 100 images from the NIPS 2017 Adversarial Attacks and Defenses Competition dataset 5.1 and ask the model to caption the image.
  • Image Classification: We randomly sample 500 images from 10 classes in the ImageNet dataset [6] and ask the model to classify the input.
  • Visual Question Answering (VQA): We utilize 100 image-question pairs from the ScienceQA dataset [26].
footnotetext: https://nips.cc/Conferences/2017/CompetitionTrack
Meanwhile, we also evaluate our FS-PSM on the original CLIP-L14 [33] on image classification for the generality.
Table 3: Experimental results on adversarial robustness of different defense methods and MLLMs on image classification tasks. The overall best results are shown in bold, and the best results without smoothing are underlined.
Model Method M-Attack [23] FOA [14] AttackVLM [51]
[…]
LLaVA- 1.5-7B Org. 0.427 8.2% 78.2% 0.437 3.8 % 81.0 % 0.458 6.0% 78.4%
Smoothed org. 0.681 90.4% 0.6% 0.697 88.8% 0.4% 0.700 88.4% 0.2%
FARE [37] 0.574 55.4% 14.6% 0.521 24.8% 39.8% 0.508 25.8% 45.0%
Smoothed FARE 0.859 75.6% 5.2% 0.712 72.0% 5.6% 0.699 74.0% 6.2%
TeCoA [29] 0.731 67.4% 2% 0.626 29.0% 20.0% 0.592 33.6% 17.8%
Smoothed TeCoA 0.819 74.4% 0.4% 0.727 57.2% 3.0% 0.711 57.4% 2.0%
CLIP -L14 Org. 0.427 18.0% 37.2% 0.437 10.0 % 52.4 % 0.458 16.0% 48.4%
Smoothed org. 0.681 92.4% 0.4% 0.697 92.0% 0.4% 0.700 92.0% 0.4%
FARE [37] 0.574 70.4% 5.2% 0.521 42.0% 14.4% 0.508 35.6% 24.4%
Smoothed FARE 0.859 85.2% 0.8% 0.712 81.6% 0% 0.699 82.4% 0%
TeCoA [29] 0.731 81.2% 2.4% 0.626 39.2% 20.2% 0.592 35.8% 24.4%
Smoothed TeCoA 0.819 88.4% 1.6% 0.727 76.0% 4.8% 0.711 73.2% 6.0%
The threat model. To comprehensively assess the robustness under strong adversaries, we employ three SOTA adversarial attacks specifically designed for MLLMs, named AttackVLM [51], M-Attack [23], and FOA [14] using the white-box setting. All attacks are implemented following their original best configurations, with the adversarial perturbation budget $\epsilon$ set to $\|\epsilon\|_{\infty}=16/255$. We also consider a stronger bound with $\|\epsilon\|_{\infty}=32/255$ for a stress test.
Compared defense methods. As this work pioneers the research for MLLMs’ certified robustness, we primarily compare our defense with adversarial training based approaches, namely FARE [37] and TeCoA [29]. Both FARE and TeCoA adopt adversarial training to obtain robust feature encoders that can be directly integrated into models such as LLaVA-1.5-7B and OpenFlamingo-9B. To ensure a fair and consistent comparison, we obtain all adversarially trained feature encoders from their official repositories, without any modification.
Implementation details. For practical inference efficiency, we set the number of samples to $n_{0}=4$ for smoothing, which provides a favorable trade-off between robustness and runtime. We train three independent PSM modules, utilizing the vanilla feature encoders from models including LLaVA-1.5-7B, OpenFlamingo-9B, and CLIP-L14. To further assess cross-model generalization, we evaluate the PSM trained on a vanilla encoder by directly applying it to adversarially trained encoders FARE and TeCoA. For all feature encoders, we set hyperparameters $\lambda_{1},~\lambda_{2},~\lambda_{4}=0.25$, $\lambda_{3}=100$, and $\sigma=0.25$. In our tables, the term ”smoothed” denotes the process of smoothing the encoder via FS and enhancing it with PSM.
Evaluation metrics. We mainly report: the Feature Cosine Similarity (FCS), the Accuracy (ACC), and the Attack Success Rate (ASR). Specifically, FCS measures the cosine similarity between clean and adversarial features extracted by the same feature encoder, reflecting its feature-space robustness. ACC denotes the proportion of correctly completed tasks, while ASR indicates the proportion of cases where the model is successfully manipulated to produce the adversarially targeted outputs. For image classification and VQA tasks, ACC and ASR are determined by whether the MLLM outputs match the correct or adversarial targets.
For image captioning, following [23], we adopt the LLM-as-a-judge protocol [52] to evaluate ACC and ASR. Specifically, we first generate the clean and adversarially targeted captions by feeding the clean and targeted inputs into the vanilla MLLM (e.g., the original LLaVA). We then obtain the predicted caption by feeding the adversarially perturbed input into the tested model (e.g., Smoothed LLaVA). The textual similarity is computed using GPTScore [9], where a task is considered successful if the GPTScore between the predicted and clean captions is $\mathbf{\geq 0.5}$, and an attack is considered successful if the GPTScore between the predicted and adversarially targeted captions is $\mathbf{\geq 0.3}$.
Table 4: Experimental results on adversarial robustness of different defense methods and MLLMs on VQA tasks. The overall best results are shown in bold, and the best results without smoothing are underlined.
Model Method M-Attack [23] FOA [14] AttackVLM [51]
[…]
LLaVA- 1.5-7B Org. 0.398 31% 28% 0.383 22% 22% 0.474 25% 27%
Smoothed org. 0.732 47% 0% 0.676 43% 0% 0.765 43% 1%
FARE [37] 0.657 47% 5% 0.590 32% 0% 0.550 38% 7%
Smoothed FARE 0.859 48% 1% 0.768 38% 0% 0.818 39% 0%
TeCoA [29] 0.788 31% 2% 0.748 31% 2% 0.667 31% 1%
Smoothed TeCoA 0.917 32% 1% 0.858 34% 1% 0.862 29% 1%
5.2 Experimental results on different tasks
Image captioning. The comparative results of different defense methods and MLLMs on the image captioning task are presented in Table 1. In this task, the adversaries aim to manipulate the MLLMs into producing the caption corresponding to a completely unrelated and maliciously chosen target image. These results demonstrate that converting MLLMs into their smoothed variants via the proposed FS-PSM yields consistently strong robustness across diverse adversarial attacks, while empirical defenses exhibit a noticeable performance drop under stronger attacks. When adversaries vary from M-Attack to FOA, the accuracy of LLaVA with FARE drops from 44% to 19%, and that of TeCoA declines from 51% to 21%. Notably, utilizing FS-PSM achieves remarkable improvements, with the accuracy of LLaVA increasing from 1% to 87% and the ASR dropping from 94% to 1% under the strongest FOA attack. Furthermore, the results demonstrate that FS-PSM possesses strong cross-model generalization; directly integrating it into FARE and TeCoA without any fine-tuning consistently enhances their robustness.
Image classification. The comparative results of different defense methods on the image classification are presented in Table 3. The adversarial objective is to mislead the model into classifying an adversarial image into a maliciously targeted class that is semantically unrelated to the original one. For image classification with CLIP-L14, we train a one-layer classification head on top of its extracted features. The results indicate that FS-PSM provides consistent and substantial robustness improvements for classification tasks. Meanwhile, those adversarial training-based defenses exhibit a noticeable performance drop under strong FOA attacks. The accuracy of FARE with LLaVA drops from 55.4% to 24.8%, and that of TeCoA declines from 67.4% to 29.0% when adversaries vary from M-Attack to FOA. Integrating FS-PSM into the vanilla model achieves remarkable improvements, with the accuracy of LLaVA increasing from 3.8% to 88.0% and the ASR dropping from 81% to 0.4% under the strongest FOA attack. Meanwhile, directly integrating the FS-PSM brings huge robustness gains for those adversarially trained feature encoders.
VQA. The comparative results of different defense methods on LLaVA-1.5-7B on the VQA are presented in Table 4. The performances of OpenFlamingo and CLIP are omitted, as these models are not directly applicable to the VQA task. In this setting, the adversarial objective is to mislead the model into selecting a wrong option, ”None of the above”, for each image–question pair. This task presents a greater challenge for pure vision-based adversaries than the previous two, as MLLMs can often infer correct answers directly from textual cues without relying heavily on visual inputs. The results demonstrate that incorporating our method yields substantial performance improvements across all attack types, significantly enhancing prediction accuracy while reducing the attack success rate to nearly zero. More implementation details for the main results are presented in our supplementary material, section S.3.
5.3 Further analysis on the main results
Reasons for FARE and TeCoA performing poorly with high FCS. Although those encoders preserve high FCS with respect to their clean inputs, the adversarial training process inherently induces a distributional shift in the feature representations. Subsequent adversarial perturbations further amplify this shift, leading to more pronounced performance degradation. In addition, the limited diversity of adversarial training data restricts their feature generalization ability, amplifying the performance drop when evaluated on downstream tasks with unseen or mismatched data distributions.
Advantages of our proposed FS with PSM. The experimental results highlight three key advantages: (1) Effectiveness: utilizing the proposed FS-PSM substantially enhances the performance of various MLLMs under adversarial attacks; (2) Trustworthiness: compared with adversarial training, FS-PSM provides more stable and consistently stronger protection across a wide range of attacks; (3) Generality: FS-PSM can be combined with other defenses to enhance the effectiveness without additional fine-tuning.
5.4 Ablation study
Table 5: The ablation study. We report the average FCSB and the average $\mathcal{R}$ for $\mathrm{FCSB}\geq 0.5$ for certified robustness. We report the average accuracy and ASR for empirical robustness under FOA attack. The clean accuracy without attack is 92.4%.
Encoder $\sigma$ Avg. FCSB at different $\epsilon$ Avg. $\mathcal{R}$ Acc ASR
0.25 0.50 0.75
CLIP-B16 / / / / / 1.6% 95.6%
FS CLIP-B16 0.50 0.586 0.253 -0.132 0.33 42.4% 0.8%
[…]
To rigorously assess the contribution of the proposed FS and each module of PSM, we adopt CLIP-B16 as the base feature encoder and evaluate the performance on the image classification task. We report both the theoretically certified robustness and the empirical performance under the strong FOA attack. The results are summarized in Table 5, where the first row indicates the performance of the vanilla CLIP-B16 model without utilizing the FS. The next three rows are the results of the smoothed CLIP-B16 with FS, smoothed CLIP-B16 with FS and mapper $\mathcal{M}$, and smoothed CLIP-B16 with FS and PSM. The results demonstrate that enforcing smoothness in the feature space substantially enhances practical robustness against adversarial perturbations, raising the accuracy from 1.6% to 91.6% and reducing the ASR from 95.6% to 0.4%. Meanwhile, each component makes a significant contribution to improving both the certified radius $\mathcal{R}$ and adversarial robustness.
More experimental results on: 1) the performance of using different purifier (a lightweight U-Net denoiser), 2) the performance under a stronger adversarial bound $\|\epsilon\|_{\infty}=32$, and 3) efficiency–robustness trade-off under various $n_{0}$, are provided in our supplementary material, section S.3.
6 Conclusion.
This work pioneers the research on establishing the certified robustness of MLLMs via a feature-space perspective. By introducing the Feature-space Smoothing (FS) framework, we show how to transform any feature encoder into a smoothed version that is equipped with a theoretical lower bound on the cosine similarity between clean and adversarial representations. Moreover, we introduce the Purifier and Smoothness Mapper (PSM), a plug-and-play module that can be seamlessly integrated into existing MLLMs to amplify the certified robustness under FS without any fine-tuning. Extensive experiments demonstrate that the proposed FS-PSM not only provides strong theoretical robustness guarantees but also achieves superior empirical protection against various attacks.
Supplementary Material
S.7 Proof of core theorem
S.7.1 Proof of lemma1
Note that:
$\nabla\Phi^{-1}(\hat{S}_{x_{c}}(\bm{x}))=\frac{\nabla\hat{S}_{x_{c}}(\bm{x})}{\Phi^{\prime}\bigr(\Phi^{-1}(\hat{S}_{x_{c}}(\bm{x}))\bigr)}.$ (S.16)
Denote $p=\hat{S}_{{\bm{x}}_{c}}({\bm{x}})$, we can get that:
$\Phi^{\prime}\bigr(\Phi^{-1}(\hat{S}_{x_{c}}(\bm{x}))\bigr)=\frac{1}{\sqrt{2\pi}}\exp\!\bigl(-\tfrac{1}{2}(\Phi^{-1}(p))^{2}\bigr)$ (S.17)
Thus, we need to prove that for any unit direction $u$.
$u\cdot\nabla\hat{S}_{{\bm{x}}_{c}}(\bm{x})\leq\frac{1}{\sqrt{2\pi}}\exp\!\bigl(-\tfrac{1}{2}(\Phi^{-1}(p))^{2}\bigr).$ (S.18)
For the left-hand side, we can get:
$\nabla\hat{S}_{x_{c}}(\bm{x})=\frac{1}{(2\pi)^{n/2}}\int_{\mathbb{R}^{n}}S_{{\bm{x}}_{c}}(\bm{t})\,({\bm{x}}-{\bm{t}})\,\exp\!\bigl(-\tfrac{1}{2}\|{\bm{x}}-{\bm{t}}\|^{2}\bigr)\,\mathrm{d}{\bm{t}},$ (S.19)
which can also be rewritten as follows:
$\mathop{\mathbb{E}}\limits_{X\sim\mathcal{N}(0,I)}\bigl[{S}_{x_{c}}(x+X)X\cdot u\bigr].$
We can now claim that the supremum of the above quantity over all encoders $f:{\bm{x}}\to{\bm{z}}$, subject to the constraint that $p=\mathbb{E}[{S}_{x_{c}}(x+X)]$, is equal to:
$\mathbb{E}\bigl[(X\cdot u)\,1\{X\cdot u\geq-\Phi^{-1}(p)\}\bigr]=\frac{1}{\sqrt{2\pi}}\exp\!\Bigl(-\tfrac{1}{2}\bigl(\Phi^{-1}(p)\bigr)^{2}\Bigr),$ (S.20)
which concludes the proof of Lemma 1.
S.7.2 Proof of Theorem1
From Lemma 1, we can get:
$\left\|{\bm{x}}-{\bm{x}}^{\prime}\right\|\geq\Phi^{-1}(\hat{S}_{{\bm{x}}}({\bm{x}}))-\Phi^{-1}(\hat{S}_{{\bm{x}}}({\bm{x}}^{\prime})).$ (S.21)
As $\left\|{\bm{x}}-{\bm{x}}^{\prime}\right\|\leq\epsilon$, we can get:
$\Phi^{-1}(\hat{S}_{{\bm{x}}}({\bm{x}}^{\prime}))\geq\Phi^{-1}(\hat{S}_{{\bm{x}}}({\bm{x}}))-\epsilon.$ (S.22)
Meanwhile, as $f_{e}({\bm{x}})$ is a unit vector, using Equation 5 in the main paper, we can get that:
$\hat{S}({\bm{x}})$ $=\tfrac{1}{2}\Big(1+\big\langle\mathbb{E}[f_{e}({\bm{x}}+\varepsilon)],\,f_{e}({\bm{x}})\big\rangle\Big)$
$=\tfrac{1}{2}\Big(1+\langle\hat{f}_{e}({\bm{x}}),\,f_{e}({\bm{x}})\rangle\Big).$ (S.23)
where $\langle\cdot,\cdot\rangle$ is the inner product. Hence, we can get:
$\langle\hat{f}_{e}({\bm{x}}),\,f_{e}({\bm{x}})\rangle=2\hat{S}({\bm{x}})-1.$ (S.24)
As $\left\|\hat{f}_{e}({\bm{x}}^{\prime})\right\|\leq 1$ and $\left\|f_{e}({\bm{x}})\right\|=1$, we can get:
${Cos}\big(\hat{f}({\bm{z}}),\,f_{e}({\bm{x}})\big)$ $=\frac{\langle\hat{f}({\bm{z}}),f_{e}({\bm{x}})\rangle}{\|\hat{f}({\bm{z}})\|}\;$ (S.25)
$\geq\;\langle\hat{f}({\bm{z}}),f_{e}({\bm{x}})\rangle$
By the monotonicity of $\Phi$,
$\hat{S}_{{\bm{x}}}({\bm{x}}^{\prime})\;\geq\;\Phi\!\Big(\Phi^{-1}\!\big(\hat{S}({\bm{x}})\big)-\epsilon\Big).$ (S.26)
Combining Equations S.24S.25S.26, we thereby can derive that:
$\mathrm{Cos}\big(\hat{f}({\bm{x}}^{\prime}),\,f_{e}({\bm{x}})\big)\;\geq\;2\,\Phi\!\Big(\Phi^{-1}\!\big(\hat{S}({\bm{x}})\big)-\epsilon\Big)-1,$ (S.27)
which concludes the proof. For noise variance $\sigma^{2}$, the same proof gives the bound with $\epsilon$ replaced by $\epsilon/\sigma$.
S.8 Training details of PSM
S.8.1 Details of fine-tuning noise purifier
The Purifier $\mathcal{P}$ is optimized in a self-supervised manner using the objective $\mathcal{L}_{p}$ defined in Equation 9. We employ the ImageNet-pretrained $256\times 256$ unconditional diffusion modelS.8.1 as the backbone of $\mathcal{P}$. All input images are resized to $256\times 256$.
To perform denoising, we instantiate a diffusion process consisting of 1,000 timesteps with a linear noise scheduler. Following common practices, the rescaled learned noise variance (i.e., ”rescale_learned_sigma”) is disabled to prevent instability during fine-tuning. We fine-tune the model for a total of 12,000 optimization steps using the AdamW optimizer, with an initial learning rate of $5\times 10^{-5}$. An exponential moving average (EMA) with a decay rate of 0.99 is maintained throughout training to stabilize updates and improve the effective generalization of the learned purifier.
footnotetext: https://github.com/openai/guided-diffusion?tab=readme-ov-file
Figure S.3: Structure of the residual smoothness mapper.
S.8.2 Details of training residual mapper
The structure of the proposed residual smoothness mapper $\mathcal{M}$ is illustrated in Figure S.3. Each block is designed to refine the purified feature representation while preserving its statistical structure and ensuring that the refinement magnitude adapts smoothly to the injected noise magnitude $\sigma$. A single block consists of: (1) noise-aware LayerNorm-FiLM($\sigma$) module, (2) multi-head attention, (3) a channel-wise MLP branch, (4) a depthwise convolution, and (5) a noise-aware residual module.
LayerNorm and FiLM($\sigma$). Each block begins by normalizing the input feature and injecting noise-level conditioning:
$\tilde{{\bm{z}}}_{i}=\mathrm{LN}(\tilde{{\bm{z}}}_{i}),\hskip 28.80008pt[\gamma(\sigma),\beta(\sigma)]=\mathrm{FiLM}(\sigma),$ (S.28)
where the FiLM module takes the scalar noise strength $\sigma$ and outputs channel-wise affine parameters through a lightweight MLP:
$\mathrm{FiLM}(\sigma)=W_{2}\,\phi(W_{1}\sigma)\in\mathbb{R}^{2D},$ (S.29)
with $\phi$ denoting a GELU activation. This modulation injects explicit noise awareness into each block: the transformation smoothly diminishes as $\sigma\to 0$, ensuring that the mapper leaves clean features nearly unchanged.
Main computation structure. Each $m_{i}(\cdot)$ contains three parallel refinement pathways:
  • Multi-head attention: Applied only in the first block, the lightweight attention layer captures long-range structural dependencies, producing global-context features $h_{\text{attn}}$.
  • MLP branch: A two-layer feed-forward network with GELU activation produces channel-wise refinement $h_{\text{mlp}}$.
  • Depthwise convolution: A depthwise 1D convolution captures local continuity in the feature sequence and contributes $h_{\text{conv}}$.
The outputs are fused as:
$h_{i}=h_{\text{attn}}+h_{\text{mlp}}+0.5\,h_{\text{conv}},$ (S.30)
where the conv branch is down-weighted following design practices for local/global feature fusion.
Noise-aware residual. We first decompose $h_{i}$ into a unit direction vector by $v_{i}={h_{i}}\setminus{\|h_{i}\|}.$ To ensure that the refinement adapts smoothly to the noise level, we modulate its magnitude through two learnable functions of $\sigma$:
$\alpha_{i}=\mathrm{softplus}\!\left(\mathrm{MLP}_{\alpha}([\sigma,\log\sigma])\right),$
$\sigma^{\beta_{i}}=\sigma^{\mathrm{softplus}(\mathrm{MLP}_{\beta}(\sigma))}.$ (S.31)
Here, $\alpha_{i}$ controls the amplitude of the refinement, while $\beta_{i}$ modulates the exponent of the noise term, enabling nonlinear noise–feature interactions. The residual update is then computed as
$\Delta{\bm{z}}_{i}=\left[(1+\gamma(\sigma))\,\bigl(v_{i}\cdot\alpha_{i}\sigma^{\beta_{i}}\bigr)+\beta(\sigma)\right]\odot\mathrm{Scale}_{i},$ (S.32)
where $\gamma(\sigma)$ and $\beta(\sigma)$ are FiLM-generated per-channel affine parameters, and $\mathrm{Scale}_{i}$ is a learnable per-channel damping factor initialized to $5\times 10^{-4}$ for stability. The next feature is then updated via
$\tilde{{\bm{z}}}_{i+1}=\tilde{{\bm{z}}}_{i}+\Delta{\bm{z}}_{i}.$ (S.33)
Overall effect of the mapper. The residual smoothness mapper thus provides the following benefits:
[…]
  • Multi-scale feature enhancement: Attention, MLP, and depth-wise convolution jointly capture global structure, per-channel adaptation, and local smoothness.
  • Stable and expressive residual learning: Fixup-style initialization and $\mathrm{Scale}_{i}$ ensure training stability even with multiple stacked residual blocks.
Collectively, $\mathcal{M}$ significantly enhances the smoothness and robustness of purified representations while avoiding distributional drift.
Table S.6: Experimental results on adversarial robustness of different defense methods and MLLMs on image classification tasks. The Mapper $\mathcal{M}$ is a lightweight U-Net. The overall best results are shown in bold, and the best results without smoothing are underlined.
Model Method M-Attack [23] FOA [14] AttackVLM [51]
[…]
LLaVA- 1.5-7B Org. 0.427 8.2% 78.2% 0.437 3.8 % 81.0 % 0.458 6.0% 78.4%
Smoothed org. 0.590 84.8% 0.2% 0.600 87.2% 0.2% 0.605 87.0% 0.2%
FARE [37] 0.574 55.4% 14.6% 0.521 24.8% 39.8% 0.508 25.8% 45.0%
Smoothed FARE 0.695 75.6% 5.2% 0.681 66.2% 1% 0.671 68.0% 1%
TeCoA [29] 0.731 67.4% 2% 0.626 29.0% 20.0% 0.592 33.6% 17.8%
Smoothed TeCoA 0.790 70.6% 0.6% 0.714 58.8% 0.8% 0.712 56.4% 0.4%
CLIP -L14 Org. 0.427 18.0% 37.2% 0.437 10.0 % 52.4 % 0.458 16.0% 48.4%
Smoothed org. 0.590 88.0% 0.4% 0.600 90.0% 0.4% 0.604 90.4% 0.4%
FARE [37] 0.574 70.4% 5.2% 0.521 42.0% 14.4% 0.508 35.6% 24.4%
Smoothed FARE 0.695 84.4% 0% 0.682 78.4% 0% 0.671 76.4% 0%
TeCoA [29] 0.731 81.2% 2.4% 0.626 39.2% 20.2% 0.592 35.8% 24.4%
Smoothed TeCoA 0.790 87.6% 1.2% 0.714 75.6% 2.4% 0.712 73.2% 2.0%
Figure S.4: Illustration of the evaluation process.
Training details. The mapper $\mathcal{M}$ is trained in a self-supervised manner using the objective $\mathcal{L}_{m}$ defined in Equation 13. We train $\mathcal{M}$ for 8 epochs with a batch size of 16. The training process employs the AdamW optimizer with an initial learning rate of $2\times 10^{-4}$ and a cosine annealing learning rate schedule.
S.9 More experimental results
S.9.1 Implementation details for main results
Image captioning. For this task, we prompt the MLLM with “Describe this image in one concise sentence, no longer than 20 words.” to generate a caption for each input image. The predicted caption is then compared against the ground-truth caption and the adversarial target caption for evaluation.
Image classification. For this task, we prompt the MLLM with ”You are a precise visual classifier. What are the main objects in this image?: 0. rooster, 1. gibbon, 2. golden_retriever, 3. goldfish, 4. hen, 5. hognose_snake, 6. ice_bear, 7. killer_whale, 8. king_crab, 9. kite, 10. shark. Output only a single integer between 0 and 10 with no explanation, no text, no punctuation.” to obtain the predicted class for each input image. All images in the dataset belong to classes 0–9, while class 10 (shark) is used as the adversarial target class. We then compute the ACC and ASR based on the predicted labels, the ground-truth labels, and the adversarial target label.
VQA. For this task, we prompt the MLLM using the provided question associated with each input image and ask it to predict the corresponding answer. An example prompt is as follows: ”You are a knowledgeable multimodal assistant. Please answer the multiple-choice question ONLY by outputting the single letter of the correct option. Do not include explanations or extra text. Question: Which ocean is highlighted? Options: A. the Arctic Ocean B. the Atlantic Ocean C. the Pacific Ocean D. the Southern Ocean E. None of the above. If the visual information is not clear or uncorrelated with the questions, you should select the choice: None of the above. Answer with exactly one letter from the options above (e.g., A).”. In this evaluation, each question has a unique correct answer among the listed options, except for “None of the above,” which is reserved for the adversarial target. The adversarial objective is to induce the MLLM to incorrectly choose “None of the above”.
An illustration of the implementation process is shown in Figure S.4.
S.9.2 Experimental results of using a lightweight U-net as noise purifier
The comparative results of different defense methods for image classification are summarized in Table S.6. In this experiment, the original diffusion-based noise purifier is replaced with a lightweight U-Net denoiser containing approximately 43.7M parameters. The results show that with this compact denoiser, FS-PSM consistently achieves strong and superior robustness, outperforming adversarial training–based baselines by a substantial margin. Furthermore, integrating FS-PSM with existing defenses yields additional robustness gains, leading to consistently favorable improvements across all evaluated settings. These results collectively demonstrate the effectiveness and generalizability of the proposed method.
S.9.3 Experimental results on large attack bound
Table S.7: Experimental results on adversarial robustness of different defense methods and MLLMs on image classification tasks. The attack bound $\epsilon=32/255$.
Model Method FOA [14]
[…]
LLaVA- 1.5-7B Org. 0.37 3.6% 78.6%
Smoothed org. 0.512 35.2% 24.2%
FARE [37] 0.408 10.4% 64.2%
Smoothed FARE 0.605 48.4% 19.2%
TeCoA [29] 0.454 12.0% 29.6%
Smoothed TeCoA 0.576 32.0% 9.6%
The experimental results under a large adversarial perturbation of $\varepsilon=32/255$ are presented in Table S.7. In addition, Figure S.5 provides visualizations of the adversarial examples generated by different attacks and perturbation bounds. These results show that nearly all existing defenses show great performance drop under such a strong attack. Under this challenging setting, the best accuracy is obtained by integrating FS-PSM with FARE, reaching 48.4% on LLaVA, while the lowest ASR (9.6%) is achieved by combining FS-PSM with TeCoA. These findings further highlight the effectiveness and adaptability of the proposed framework, even against large-magnitude adversarial perturbations.
Figure S.5: Visualization of the adversarial examples.
S.9.4 Analysis of robustness-efficiency trade-off
Table S.8: Analysis of the robustness-efficiency trade-off of using different $n_{0}$.
Model $n_{0}$ Robustness Efficiency
[…]
LLaVA- 1.5-7B (diffusion purifier) 1 0.595 86.2% 0.6% 0.28
2 0.652 88.4% 0.4% 0.41
4 0.697 88.8% 0.4% 0.66
8 0.719 89.6% 0.4% 1.12
16 0.728 89.2% 0.4% 2.10
LLaVA- 1.5-7B (lightweight purifier) 1 0.494 84.4% 0.2% 0.20
2 0.565 85.4% 0% 0.24
4 0.600 87.2% 0.2% 0.32
8 0.623 87.8% 0.4% 0.49
16 0.643 86.8% 0.4% 0.84
The robustness–efficiency trade-off results are summarized in Table S.8. In this experiment, we adopt the strong FOA attack as the adversary and report the average inference time per sample on LLaVA-1.5-7B using a single RTX 4090 GPU. The results show that increasing the sampling number $n_{0}$ consistently improves the FCS under adversarial perturbations. Specifically, increasing $n_{0}$ from 0 to 16 raises FCS from approximately 0.595 to 0.728, which aligns with our theoretical analysis in Section 3. However, this robustness gain comes at a substantial computational cost. For example, increasing $n_{0}$ from 4 to 16 raises the inference latency from 0.66 seconds to 2.10 seconds per sample, revealing an inherent trade-off between robustness and efficiency.

### v2

@@
-Provable Robustness in Multimodal Large Language Models via Feature Space Smoothing
+Feature-Space Adversarial Robustness Certification for Multimodal Large Language Models
 Abstract
 Multimodal large language models (MLLMs) exhibit strong capabilities across diverse applications, yet remain vulnerable to adversarial perturbations that distort their feature representations and induce erroneous predictions.
-To address this vulnerability, we propose the Feature-space Smoothing (FS) and theoretically prove that FS offers certified robustness on the feature representations of MLLMs.
-Specifically, FS transforms any feature encoder into a smoothed variant that is guaranteed to maintain a certified lower bound on the feature cosine similarity between clean and adversarial representations under $\ell_{2}$-bounded attacks.
-Moreover, we indicate that the value of this Feature Cosine Similarity Bound (FCSB) derived from FS can be improved by enlarging the defined Gaussian robustness score on the vanilla encoder.
-Building upon this, we introduce the Purifier and Smoothness Mapper (PSM), a plug-and-play module that improves the Gaussian robustness score of MLLMs and thus enhances their certified robustness under FS, without requiring any retraining on MLLMs.
-We demonstrate that the FS with PSM not only provides a strong theoretical robustness guarantee but also exhibits superior empirical performance compared to adversarial training.
-Extensive experiments across diverse MLLMs and downstream tasks indicate the effectiveness of the FS-PSM, reducing the Attack Success Rate (ASR) of various white-box attacks from nearly 90% to about 1%.
+To address this vulnerability, we propose Feature-space Smoothing (FS), a general framework that provides certified robustness guarantees at the feature representation level of MLLMs.
+We theoretically prove that FS converts a given feature extractor into a smoothed variant that is guaranteed a certified lower bound on the cosine similarity between clean and adversarial features under $\ell_{2}$-bounded perturbations.
+Moreover, we establish that the value of this Feature Cosine Similarity Bound (FCSB) is determined by the intrinsic Gaussian robustness score of the given encoder.
+Building on this insight, we introduce the Gaussian Smoothness Booster (GSB), a plug-and-play module that enhances the Gaussian robustness score of pretrained MLLMs, thereby strengthening the robustness guaranteed by FS, without requiring additional MLLM retraining.
+Extensive experiments demonstrate that applying the FS to various MLLMs yields strong certified feature-space robustness and consistently leads to robust task-oriented performance across diverse applications.
+Machine Learning, ICML
 1 Introduction
-The emergence of the Multimodal Large Language Models (MLLMs), such as GPT-5, Gemini 2.5, and Claude 3.7 Sonnet, has fundamentally reshaped existing working paradigms and significantly advanced societal productivity.
-Despite their remarkable capabilities across a broad spectrum of real-world tasks, these models still encounter critical safety challenges, such as adversarial vulnerabilities [51, 5, 23, 14].
-Figure 1: Illustration of the FS-PSM, which guarantees that the cosine similarity of the adversarial and clean features extracted by MLLM’s encoder is larger than FCSB for robust predictions.
-Adversaries can manipulate predictions of the MLLMs to a malicious state by injecting subtle and imperceptible perturbations to the clean inputs, exploiting the models’ insufficient local smoothness and uncontrolled Lipschitz continuity [10, 4, 12, 47].
+The emergence of the Multimodal Large Language Models (MLLMs), such as GPT-5 (OpenAI, 2025), Gemini 3 Pro (DeepMind, 2025), and Claude Sonnet 4.5 (Anthropic, 2025), has fundamentally reshaped existing working paradigms and significantly advanced societal productivity.
+Despite their remarkable capabilities across a broad spectrum of real-world tasks, these models still encounter critical safety challenges, such as adversarial vulnerabilities (Zhao et al., 2023; Cui et al., 2024; Li et al., 2025b; Jia et al., 2025).
+Figure 1: Illustration of the FS-GSB, which guarantees that the cosine similarity of the adversarial and clean features extracted by MLLM’s encoder is larger than FCSB for robust predictions.
+Adversaries can manipulate predictions of the MLLMs to a malicious state by injecting subtle and imperceptible perturbations to the clean inputs, exploiting the models’ insufficient local smoothness and uncontrolled Lipschitz continuity (Goodfellow et al., 2014; Cohen et al., 2019; Hein and Andriushchenko, 2017; Xia et al., 2024b).
 Countermeasures towards those threats can be roughly classified into empirical defense and certified defense.
-Typical empirical approaches include adversarial training [27, 35, 41, 45, 2, 37, 28, 29] and input purification [27, 35, 41, 45, 2, 37, 28, 29].
-Despite their demonstrated empirical effectiveness, the multimodal nature of MLLMs also poses a great challenge for existing adversarial defense methods.
-Since these models accept heterogeneous inputs across diverse domains, ensure a robust encoder via adversarial training that can generalize to various scenarios is challenging and computationally expensive.
-Moreover, these approaches lack formal robustness guarantees and remain susceptible to stronger or adaptive adversaries [38, 3, 44].
-In contrast, certified approaches aim to guarantee that the model returns a constant prediction result within a certain range, usually a $\ell_{2}$ or $\ell_{\infty}$-norm constrained area [34, 43, 11, 15, 4, 47, 36].
-However, most previous certified defense approaches [4, 47, 42, 11, 36] predominantly assume the prediction of one-dimensional outputs (e.g., class label), thereby limiting their applicability to more general tasks on MLLMs.
-To address the aforementioned challenges, we propose the Feature-space Smoothing (FS), a provable defense method that offers certified robustness for the feature representations of MLLMs.
-FS smooths the vanilla feature encoder of the MLLM, and the resulted smoothed encoder is guaranteed to maintain a provable lower bound on the cosine similarity between clean and adversarial representations under $\ell_{2}$-norm bounded attacks.
-Moreover, we demonstrate that the Feature Cosine Similarity Bound (FCSB) of the smoothed feature encoder is intrinsically determined by a Gaussian robustness score, which is defined to measure the prediction consistency of the vanilla feature extractor under Gaussian noise.
-Nonetheless, without Gaussian noise–augmented training, the Gaussian robustness of these MLLMs remains limited, leading to a correspondingly suboptimal FCSB.
-To address this and avoid the heavy computational cost of fine-tuning MLLMs, we propose a plug-and-play Purifier and Smoothness Mapper (PSM) that effectively enhance the Gaussian robustness score for the vanilla feature encoder in a training-free way.
-Specifically, the purifier operates prior to feature extraction and is trained to denoise Gaussian perturbations and maximize the Gaussian robustness score.
-Meanwhile, the smoothness mapper performs post-extraction refinement and is trained to preserve the feature distribution while further enhancing this score without fine-tuning the encoder.
-These two components of PSM work synergistically to enhance the certified robustness of MLLMs under FS.
-To implement our PSM, we employ a pre-trained guided diffusion model as the Gaussian noise purifier and design a noise-aware residual network as the smoothness mapper.
-The PSM is optimized using the proposed utility–robustness loss and trained on data from diverse visual domains to enhance the Gaussian robustness while preserving the feature utility for the encoder.
-To comprehensively assess the performance, we evaluate our method against state-of-the-art (SOTA) adversarial attacks tailored for MLLMs under the white-box setting and compare it with strong adversarial training defense.
-Experimental results demonstrate that our FS with PSM enhanced provides strong protection for various MLLMs under diverse downstream tasks.
-Overall, the main contributions of this work are:
-  • We propose the Feature-space Smoothing (FS) to turn any feature encoder of MLLMs into a smoothed version, and theoretically prove that the smoothed encoder maintains a certified lower bound on the feature cosine similarity between clean and adversarial representations.
-  • We propose a plug-and-play Purifier and Smoothness Mapper (PSM) that effectively enhances the value of the FCSB of the smoothed encoder in a training-free manner.
-  • We conduct extensive experiments demonstrating that integrating our proposed FS-PSM greatly enhances the adversarial performance of various MLLMs and dramatically reduces the ASR under various white-box attacks.
+Typical empirical approaches include adversarial training (Madry et al., 2018; Rebuffi et al., 2021; Wang et al., 2024c; Xhonneux et al., 2024; Casper et al., 2024; Schlarmann et al., 2024; Malik et al., 2025; Mao et al., 2023) and input purification (Nie et al., 2022; Yoon et al., 2021; Lee and Kim, 2023; Lei et al., 2025; Zollicoffer et al., 2025).
+Despite their demonstrated empirical effectiveness, these approaches lack formal robustness guarantees and remain susceptible to stronger adversaries (Tramer et al., 2020; Chen et al., 2023; Wu et al., 2020; Olivier and Raj, 2023).
+Moreover, the multimodal nature of MLLMs also poses a great challenge for existing adversarial training methods.
+Since they accept heterogeneous inputs across diverse domains, ensure a robust encoder via adversarial training that can generalize to various scenarios is challenging and computationally costly.
+In contrast, certified approaches aim to guarantee that the model returns a constant prediction result within a certain range, usually a $\ell_{2}$ or $\ell_{\infty}$-norm constrained area (Raghunathan et al., 2018; Wong and Kolter, 2018; Hao et al., 2022; Kakizaki et al., 2023; Cohen et al., 2019; Xia et al., 2024b; Salman et al., 2019).
+However, most previous certified defense approaches (Cohen et al., 2019; Xia et al., 2024b; Wang et al., 2024b; Hao et al., 2022; Salman et al., 2019) predominantly assume the prediction of one-dimensional outputs (e.g., class label), thereby limiting their applicability to more general tasks on MLLMs.
+To address the aforementioned challenges, we propose the Feature-space Smoothing (FS), a provable defense method that offers certified robustness on the feature representations of MLLMs.
+Specifically, FS smooths the vanilla feature encoder of the MLLM, and the resulted smoothed encoder is guaranteed to maintain a provable lower bound on the cosine similarity between clean and adversarial representations under $\ell_{2}$-norm bounded attacks.
+Moreover, we demonstrate that this Feature Cosine Similarity Bound (FCSB) of the smoothed feature encoder is intrinsically determined by a defined Gaussian robustness score, which measures the prediction consistency of the vanilla feature extractor under Gaussian noise.
+Nonetheless, the Gaussian robustness of these MLLMs remains limited without Gaussian noise augmented training, leading to a correspondingly suboptimal FCSB.
+To address this whiling avoiding the high computational cost of fine-tuning MLLMs, we propose a plug-and-play Gaussian Smoothness Booster (GSB), consisting of a lightweight Gaussian denoiser and a residual smoothness mapper.
+Specifically, the denoiser is applied prior to the feature encoder and is trained to mitigates Gaussian noises and improves the resulting robustness score.
+In parallel, the smoothness mapper operates after feature extraction for feature-level refinement.
+It is trained to preserve the original feature distribution while further enhancing the Gaussian robustness score.
+These two components of GSB work synergistically to enhance the certified robustness of MLLMs provided by FS.
+The GSB is optimized using the proposed utility–robustness loss and trained on data from diverse visual domains to enhance the Gaussian robustness while preserving the feature utility for the encoder.
+To comprehensively assess the performance, we evaluate our method against state-of-the-art (SOTA) adversarial attacks tailored for MLLMs under the white-box setting and compare it with advanced adversarial training methods.
+Extensive experimental results demonstrate that our FS not only provides a strong certified robustness guarantee on the feature representation but also exhibits superior adversarial robustness for various MLLMs under diverse downstream tasks.
+Overall, the contributions of this work are:
+  • We propose the Feature-space Smoothing (FS) to turn any feature encoder into a smoothed version, and theoretically prove that the smoothed encoder maintains a certified lower bound on the feature cosine similarity between clean and adversarial representations.
+  • We propose the Gaussian Smoothness Booster (PSB) that effectively enhances the FCSB of the smoothed encoder in MLLMs via a plug-and-play manner.
+  • We conduct extensive experiments demonstrating that integrating our proposed FS-GSB greatly enhances the adversarial robustness of various MLLMs and dramatically reduces the ASR under various white-box attacks.
 2 Related Work
 Adversarial attacks on MLLMs.
-While MLLMs continue to achieve remarkable performance across diverse applications, extensive works [31, 5, 51, 14, 23, 39, 49, 50, 48] have exposed their adversarial vulnerabilities, raising serious safety concerns.
-Early work, such as AttackVLM [51], explores transferable attacks by disrupting the feature representations of CLIP [33] and BLIP [19], showing strong adversarial transferability among open-source models but limited effectiveness against closed-source commercial systems.
-More recent approaches, such as M-Attack [23] and FOA-Attack [14], further advance this direction by leveraging multi-extractor ensembles and feature-space alignment, achieving over 90% targeted attack success rates on image-captioning tasks against powerful closed-sourced commercial MLLMs (e.g., ChatGPT-4o).
+While MLLMs continue to achieve remarkable performance across diverse applications, extensive works (Qi et al., 2024; Cui et al., 2024; Zhao et al., 2023; Jia et al., 2025; Li et al., 2025b; Wang et al., 2024a; Zhang et al., 2024, 2025; Xie et al., 2025) have exposed their adversarial vulnerabilities, raising serious safety concerns.
+Early work, such as AttackVLM (Zhao et al., 2023), explores transferable attacks by disrupting the feature representations of CLIP (Radford et al., 2021) and BLIP (Li et al., 2023a), showing strong adversarial transferability among open-source models but limited effectiveness against closed-source commercial systems.
+More recent approaches, such as M-Attack (Li et al., 2025b) and FOA-Attack (Jia et al., 2025), further advance this direction by leveraging multi-extractor ensembles and feature-space alignment, achieving over 90% targeted attack success rates on image-captioning tasks against powerful closed-sourced commercial MLLMs (e.g., ChatGPT-4o).
 This highlights an urgent need for trustworthy defenses that provide effective and provable protection for MLLMs.
@@
 Adversarial defense methods can be broadly classified into empirical and provable approaches.
-Empirical defenses for MLLMs mainly include adversarial training [27, 35, 41, 45, 2, 37, 28, 29], which enhances robustness by augmenting training data with adversarial examples, and input purification [30, 17, 22], which employs generative mechanisms such as diffusion models or autoencoders to recover clean inputs prior to inference.
-Recent studies [28, 29, 37] have revealed that utilization of adversarially trained CLIP-feature encoders can enhance adversarial robustness for MLLMs.
-However, these methods lack formal robustness guarantees and remain vulnerable to adaptive and unseen threats [38, 3, 44].
+Empirical defenses for MLLMs mainly include adversarial training (Madry et al., 2018; Rebuffi et al., 2021; Wang et al., 2024c; Xhonneux et al., 2024; Casper et al., 2024; Schlarmann et al., 2024; Malik et al., 2025; Mao et al., 2023), which enhances robustness by augmenting training data with adversarial examples, and input purification (Nie et al., 2022; Lei et al., 2025; Li et al., 2025a), which employs generative mechanisms such as diffusion models or autoencoders to recover clean inputs prior to inference.
+Recent studies (Malik et al., 2025; Mao et al., 2023; Schlarmann et al., 2024) have revealed that utilization of adversarially trained CLIP-feature encoders can enhance adversarial robustness for MLLMs.
+However, these methods lack formal robustness guarantees and remain vulnerable to adaptive and unseen threats (Tramer et al., 2020; Chen et al., 2023; Wu et al., 2020).
 Moreover, adversarial training demands costly retraining and often leads to degradation in clean performance.
 Alternatively, certified defenses aim to provide mathematically provable robustness guarantees.
-The use of Gaussian smoothing for certified robustness was initially introduced for classification models [4, 18, 16], yet its theoretical formulation is restricted to one-dimensional outputs, limiting its applicability to tasks such as auto-regression or multimodal generation.
-To overcome these limitations and ensure trustworthy protection for MLLMs, we propose the Feature-space Smoothing (FS), a defense mechanism that establishes provable adversarial robustness for MLLMs.
-3 Feature Space Smoothing
+The use of Gaussian smoothing for certified robustness was initially introduced for classification models (Cohen et al., 2019; Li et al., 2018; Lecuyer et al., 2019), yet its theoretical formulation is restricted to one-dimensional outputs, limiting its applicability to tasks such as auto-regression or multimodal generation.
+To overcome these limitations and ensure trustworthy protection for MLLMs, we propose the Feature-space Smoothing (FS), a defense mechanism that establishes provable adversarial feature robustness for MLLMs.
+3 Feature-Space Smoothing
 3.1 Preliminary
@@
 For each input $\bm{x}$}, the adversarial attacks aim to find an adversarial input ${\bm{x}}^{\prime}={\bm{x}}+{\bm{\delta}}$ that misleads the model by solving:
-$\vskip-2.84526pt\mathop{\max}\limits_{{{\bm{x}}}+{\bm{\delta}}\in\mathcal{B}_{\epsilon}({\bm{x}})}\mathcal{L}\left({\mathcal{F}\left({{{\bm{x}}}}\right),\mathcal{F}\left({{{\bm{x}}}+{\bm{\delta}}}\right)}\right).\vskip-2.84526pt$ (1)
-Adversarial effects on feature representations: While adversarial attacks primarily aim to alter the model’s predictions, numerous studies [14, 46, 20, 13, 21, 23, 40, 32, 8] have shown that successful attacks typically induce substantial distortions in the model’s feature representations.
+$\vskip-2.84526pt\mathop{\max}\limits_{{{\bm{x}}}+{\bm{\delta}}\in\mathcal{B}_{\epsilon}({\bm{x}})}\mathcal{L}\left({\mathcal{F}\left({{{\bm{x}}}}\right),\mathcal{F}\left({{{\bm{x}}}+{\bm{\delta}}}\right)}\right).$ (1)
+Adversarial effects on feature representations: While adversarial attacks primarily aim to alter the model’s predictions, numerous studies (Jia et al., 2025; Xia et al., 2024a; Li et al., 2020; Huang et al., 2019; Li et al., 2023b; Ding et al., 2024) have shown that successful attacks typically induce substantial distortions in the model’s feature representations.
 Let $\bm{x}$$^{\prime}$ denote the adversarial example with adversarial feature $\bm{z}$$^{\prime}$.
@@
 Then it turns an arbitrary base classifier $\mathcal{F}$ into a smoothed version $\hat{\mathcal{F}}$ that possesses $\ell_{2}$} certified robustness guarantees.
-The smoothed classifier $\hat{\mathcal{F}}$ returns whichever the class the base classifier $\mathcal{F}$ is most likely to return among the distribution ${\bm{x}}+\bm{\varepsilon}\sim\mathcal{N}({\bm{x}},{\sigma^{2}}\bm{I})$, which is:
-$\begin{array}[]{l}\hat{\mathcal{F}}({\bm{x}})=\mathop{\arg\max}\limits_{c\in\mathcal{Y}}{\mathbb{P}}(\mathcal{F}({\bm{x}}+\bm{\varepsilon})=c).\vskip-5.69054pt\end{array}$ (2)
+The smoothed classifier $\hat{\mathcal{F}}$ returns whichever class the base classifier $\mathcal{F}$ is most likely to return among the distribution ${\bm{x}}+\bm{\varepsilon}\sim\mathcal{N}({\bm{x}},{\sigma^{2}}\bm{I})$, which is:
+$\begin{array}[]{l}\hat{\mathcal{F}}({\bm{x}})=\mathop{\arg\max}\limits_{c\in\mathcal{Y}}{\mathbb{P}}(\mathcal{F}({\bm{x}}+\bm{\varepsilon})=c).\vskip-2.84526pt\end{array}$ (2)
 RS then guarantees a certified radius $\mathcal{R}$ for this smoothed classifier $\hat{\mathcal{F}}$.
@@
 While previous RS provides effective certified protection for classification models, it suffers from two limitations.
-First, the theoretical framework of RS inherently restricts its certification to classification tasks (e.g.
-possibility of input $\bm{x}$} belongs to a certain class $c$).
+First, the theoretical framework of RS inherently restricts its certification to classification tasks (e.g.possibility of input $\bm{x}$} belongs to a certain class $c$).
 Second, estimating $\mathbb{P}(\mathcal{F}({\bm{x}}+\bm{\varepsilon})=c)$ in Equation 2, incurs substantial computational overhead, since each estimation requires multiple forward passes through the entire model.
@@
 $\bm{\varepsilon}\sim\mathcal{N}(0,I)$ denotes Gaussian noise with zero mean and standard deviation one.
-The smoothed feature encoder outputs the expectation of feature representations over the Gaussian distribution $\mathcal{N}({\bm{x}},I)$.
-Figure 2: The training framework of the PSM.
-The purifier performs pre-processing and the smoothness mapper refines post-extracted features to enhance the Gaussian robustness.
-Parameters of MLLMs are frozen, and the purifier and mapper are optimized with $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\mathcal{M}}$.
-For evaluation, the input $\bm{x}$} is replaced with adversarial input $\bm{x}$$^{\prime}$ and the forward calculation marked by orange color will be removed.
+The smoothed feature encoder $\hat{f}_{e}(\bm{x})$ outputs the expectation of feature representations over the Gaussian distribution $\mathcal{N}({\bm{x}},I)$.
+Figure 2: The training framework of the GSB.
+The denoiser performs pre-processing, and the smoothness mapper refines post-extracted features to enhance the Gaussian robustness.
+Parameters of MLLMs are frozen, and the denoiser and mapper are optimized with $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\mathcal{M}}$.
+For evaluation, $\bm{x}$} is replaced with adversarial input $\bm{x}$$^{\prime}$, and the forward calculation marked in orange color will be removed.
 Gaussian robustness score. Define ${S_{\bm{x_{t}}}}({\bm{x}})$ as the score function that evaluates feature discrepancy between an input $\bm{x}$} and a targeted example ${\bm{x}}_{t}$, which is:
-$\vskip-5.69054ptS_{{\bm{x}}_{t}}({\bm{x}})=\tfrac{1}{2}\Big(1+\mathrm{Cos}\big(f_{e}({\bm{x}}),\,f_{e}({\bm{x}}_{t})\big)\Big)$ (4)
+$\vskip-2.84526ptS_{{\bm{x}}_{t}}({\bm{x}})=\tfrac{1}{2}\Big(1+\mathrm{Cos}\big(f_{e}({\bm{x}}),\,f_{e}({\bm{x}}_{t})\big)\Big)$ (4)
 where ${Cos}(\cdot,\cdot)$ denotes the cosine similarity and $S_{{\bm{x}}_{t}}({\bm{x}})\in[0,1]$. The Gaussian robustness score $\hat{S}({\bm{x}})$ is defined as:
@@
 $=\tfrac{1}{2}\!\left(1+\mathop{\mathbb{E}}\limits_{\bm{\varepsilon}\sim\mathcal{N}(0,I)}\!\left[\mathrm{Cos}\big(f_{e}({\bm{x}}+\bm{\varepsilon}),\,f_{e}({\bm{x}})\big)\right]\right),$
-The score $\hat{S}({\bm{x}})$ evaluates the expected cosine similarity between the feature representation of a Gaussian-perturbed input ${\bm{x}}+\bm{\varepsilon}$ and that of the clean input $\bm{x}$}, characterizing the Gaussian robustness of the vanilla feature encoder ${f}_{e}({\bm{x}})$.
+The score $\hat{S}({\bm{x}})$ evaluates the expected cosine similarity between the feature representation of a Gaussian-perturbed input ${\bm{x}}+\bm{\varepsilon}$ and that of the clean input $\bm{x}$}, characterizing the Gaussian robustness of the vanilla feature encoder ${f}_{e}$.
 Notably, we prove that this score $\hat{S}({\bm{x}})$ preserves a good Lipschitz property, which serves as a theoretical foundation for proving the certified robustness for $\hat{f}_{e}({\bm{x}})$.
-Lemma 1 (Lipschitz property for the Gaussian robustness score).
+Lemma 3.1 (Lipschitz property for the Gaussian robustness score).
 Let $\Phi(a)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{a}\exp(-\frac{1}{2}s^{2})ds$ be a standard Gaussian cumulative distribution function and $\Phi^{-1}$ be its inverse.
 For any feature encoder $f_{e}:{\bm{x}}\rightarrow{\bm{z}}$, the mapping ${\bm{x}}\rightarrow\Phi^{-1}(\hat{S}(\bm{x}))$ is 1Lipschitz1-Lipschitz.
-The proof of Lemma 1 is in the supplementary material, Section S.1.
+The proof of Lemma 3.1 is in the Appendix, Section A.1.1.
 It implies that the mapping from $\bm{x}$} to $\Phi^{-1}(\hat{S}({\bm{x}}))$ exhibits strong adversarial robustness, as it satisfies a $1$-Lipschitz constraint.
@@
 The following Theorem establishes an explicit relationship between this score and the lower bound on the adversarial feature cosine similarity.
-Theorem 1 (Certified lower bound on the adversarial feature cosine similarity).
+Theorem 3.2 (Certified lower bound on the adversarial feature cosine similarity).
 For any feature encoder ${f}_{e}$ and its smoothed version $\hat{f}_{e}$, let $\bm{x}$} and $\bm{x}$$^{\prime}$ be clean and adversarial inputs with $\|{\bm{x}}^{\prime}-{\bm{x}}\|\leq\epsilon$.
 The cosine similarity between the adversarial feature $\hat{f}_{e}({\bm{x}}^{\prime})$ and clean feature ${f}_{e}({\bm{x}})$ satisfies: $\mathrm{Cos}(\hat{f}_{e}({\bm{x}}^{\prime}),f_{e}({\bm{x}}))\geq 2{\Phi}\left({\Phi^{-1}}(\hat{S}({\bm{x}}))-\epsilon\right)-1$.
-Denote $2{\Phi}\left({\Phi^{-1}}(\hat{S}({\bm{x}}))-\epsilon\right)-1$ as the Feature Cosine Similarity Bound (FCSB). Theorem 1 reveals that:
+Denote $2{\Phi}\left({\Phi^{-1}}(\hat{S}({\bm{x}}))-\epsilon\right)-1$ as the Feature Cosine Similarity Bound (FCSB). Theorem 3.2 reveals that:
   • By FS, we can turn any given feature encoder $f_{e}$ into a smoothed version $\hat{f}_{e}$ that maintains a FCSB between the adversarial and clean feature representations.
   • By maximizing the robustness score $\hat{S}({\bm{x}})$ of the given feature encoder $f_{e}$, we can effectively enhance the value of FCSB derived on its smoothed version $\hat{f}_{e}$.
-The proof of Theorem 1 is in the supplementary material, Section S.1.
-Corollary 1 (Certified radius $\mathcal{R}$ for adversarial cosine similarity $\geq$ 0.5).
+The proof of Theorem 3.2 is in the Appendix, Section A.1.2.
+Corollary 3.3 (Certified radius $\mathcal{R}$ for adversarial cosine similarity $\geq$ 0.5).
 Let $\bm{x}$} be the clean input, and $\bm{x}$$^{\prime}$ be the adversarial input.
@@
 $\vskip-2.84526pt\mathcal{R}=\Phi^{-1}\!\big(\hat{S}({\bm{x}})\big)-\Phi^{-1}(0.75).$ (6)
-Building upon Theorem 1, Corollary 1 establishes a certified radius $\mathcal{R}$ for the smoothed feature encoder.
+Building upon Theorem 3.2, Corollary 3.3 establishes a certified radius $\mathcal{R}$ for the smoothed feature encoder.
 For any adversarial perturbation satisfying $\|\delta\|_{2}\leq\mathcal{R}$, the FCSB of the smoothed encoder is guaranteed to remain above 0.5.
-Why feature-space smoothing for MLLMs? Unlike previous smoothing methods, performing smoothing in the feature space offers several advantages:
-  • Efficiency: Compared to the previous RS that smooths the entire model $\mathcal{F}$, smoothing the feature encoder $f_{e}$, which is substantially lighter than the whole MLLM $\mathcal{F}$, significantly reduces the computational time of multiple forward calculations required by smoothing.
-  • Generality: The FS provides certified robustness at the feature representation level, making it applicable to various downstream tasks (e.g.
-image captioning, classification and visual question answering).
-  • Effectiveness: Since feature representations play a critical role in the final prediction, guaranteeing a trustworthy feature representation effectively improves the prediction reliability and robustness under adversarial attacks.
-4 Purifier and Smoothness Mapper
-Theorem 1 and Corollary 1 reveal an intriguing robustness property of the smoothed feature encoder $\hat{f}_{e}$.
-However, MLLMs’ feature encoders $f_{e}$ generally exhibits limited Gaussian robustness, which restricts the value of FCSB derived on $\hat{f}_{e}$.
+4 Gaussian Smoothness Booster
+Theorem 3.2 and Corollary 3.3 reveal an intriguing robustness property of the smoothed feature encoder $\hat{f}_{e}$.
+However, feature encoders $f_{e}$ of MLLMs generally exhibit limited Gaussian robustness, which restricts the value of FCSB derived on $\hat{f}_{e}$.
 One solution is to estimate the smoothness score $\hat{S}({\bm{x}})$ through Monte Carlo sampling and training $f_{e}$ to maximize this score via gradient backpropagation.
-However, this could reduce the adaptability and practicality of our FS protection, as fine-tuning and re-aligning the LLM with multi-modal feature encoders is highly complex and costly.
-This thereby motivates us to propose the Purifier and Smoothness Mapper (PSM), a plug-and-play module that can be seamlessly integrated with MLLMs to enhance its Gaussian robustness score $\hat{S}(x)$.
-The training framework of our proposed PSM is shown in Figure 2.
-The purifier $\mathcal{P}$ operates before feature extraction to denoise the Gaussian perturbations, and the smoothness mapper $\mathcal{M}$ performs post-extraction to do the feature refinement.
+However, this could reduce the adaptability and practicality of the proposed FS framework, as fine-tuning and re-aligning the LLM is highly complex and costly.
+This thereby motivates us to propose the Gaussian Smoothness Booster (GSB), a plug-and-play module that can be seamlessly integrated with MLLMs to enhance its Gaussian robustness score $\hat{S}(x)$.
+The training framework of our proposed GSB is shown in Figure 2.
+The denoiser $\mathcal{P}$ operates before feature extraction to denoise the Gaussian perturbations, and the smoothness mapper $\mathcal{M}$ performs post-extraction to do the feature refinement.
 Those two modules work together to enhance the Gaussian robustness of the given feature encoder.
-4.1 Noise purifier
-To denoise the Gaussian noise, $\mathcal{P}$ is trained to minimize the reconstruction loss $l_{mse}$, defined as:
+4.1 Gaussian Denoiser
+To perform the Gaussian denoising, the denoiser $\mathcal{P}$ is trained to minimize the reconstruction loss $l_{mse}$, defined as:
 $\vskip-2.84526ptl_{mse}=\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}{\bm{x}}\sim\mathcal{D},~\bm{\varepsilon}\sim\mathcal{N}(0,\sigma^{2}I)\end{subarray}}\|{\bm{x}}-\mathcal{P}({\bm{x}}+\mathbf{\varepsilon})\|,$ (7)
@@
 which encourages feature consistency between the purified and clean representations.
-Practically, we adopt the ImageNet-pretrained guided-diffusion [7] as our purifier $\mathcal{P}$ and perform one-step diffusion for denoising in our main experiments.
 We then fine-tune this model on the dataset $\mathcal{D}$ using the purifier loss $\mathcal{L}_{\mathcal{P}}$, which is:
-$\vskip-2.84526pt\mathcal{L}_{\mathcal{P}}=l_{\mathrm{diff}}+\lambda_{1}l^{\mathcal{P}}_{\mathrm{rb}}+\lambda_{2}l_{mse},$ (9)
-where $l_{\mathrm{diff}}$ denotes the original diffusion loss, and $\lambda_{1}$ and $\lambda_{2}$ are the weighting coefficients.
-More details of the fine-tuning process on $\mathcal{P}$ can be found in the supplementary material, Section S.2.
+$\vskip-2.84526pt\mathcal{L}_{\mathcal{P}}=\ l^{\mathcal{P}}_{\mathrm{rb}}+\lambda_{1}l_{mse},$ (9)
+where $\lambda_{1}$ is the weighting coefficient. More details of the fine-tuning process on $\mathcal{P}$ can be found in the Appendix A.2.1.
 4.2 Residual Smoothness Mapper
@@
 The overall training loss $\mathcal{L}_{\mathcal{M}}$ is defined as:
-$\vskip-2.84526pt\mathcal{L}_{\mathcal{M}}=l^{\mathcal{M}}_{\mathrm{rb}}+\lambda_{3}l_{\mathrm{stats}}+\lambda_{4}l_{\mathrm{id}},$ (13)
-where $\lambda_{3}$ and $\lambda_{4}$ are the weighting coefficients. More details of the training process on the residual smoothness mapper can be found in the supplementary material, Section S.2.
-4.3 Further discussion on PSM
-Certified robustness for the encoder $f_{e}$ with PSM. Let $f^{\prime}_{e}$ denote the feature encoder integrated with the proposed PSM. Then the forward process can be formulated as:
+$\vskip-2.84526pt\mathcal{L}_{\mathcal{M}}=l^{\mathcal{M}}_{\mathrm{rb}}+\lambda_{2}l_{\mathrm{stats}}+\lambda_{3}l_{\mathrm{id}},$ (13)
+where $\lambda_{2}$ and $\lambda_{3}$ are the weighting coefficients. More details of the training process on the residual smoothness mapper can be found in the Appendix A.2.2.
+Algorithm 1 Training algorithm of GSB
+0:  Dataset $\mathcal{D}$, feature encoder $f_{e}$, Denoiser $\mathcal{P}$, mapper $\mathcal{M}$, sampling number $n_{0}$, noise std $\sigma$, loss weights $\lambda_{1},\lambda_{2},\lambda_{3}$.
+1:  Stage 1: Train denoiser $\mathcal{P}$.
+2:  for each batch ${\bm{x}}\sim\mathcal{D}$ do
+3:  ${\bm{z}}\leftarrow f_{e}({\bm{x}})$
+4:  for $i=1$ to $n_{0}$ do
+5:   Sample $\bm{\varepsilon}^{i}\sim\mathcal{N}(0,\sigma^{2}I)$
+6:   $\tilde{{\bm{x}}}^{i}\leftarrow\mathcal{P}({\bm{x}}+\bm{\varepsilon}^{i})$
+7:   $\tilde{{\bm{z}}}^{i}\leftarrow f_{e}(\tilde{{\bm{x}}}^{i})$
+8:  end for
+9:  Compute $l^{\mathcal{P}}_{\mathrm{rb}}$ and $l_{\mathrm{mse}}$ using Eqs. 7 and 8.
+10:  Update $\mathcal{P}$ by gradient descent on $\nabla_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}$
+11:  end for
+12:  Stage 2: Train smoothness mapper $\mathcal{M}$.
+13:  for each batch ${\bm{x}}\sim\mathcal{D}$ do
+14:  ${\bm{z}}\leftarrow f_{e}({\bm{x}})$
+15:  for $i=1$ to $n_{0}$ do
+16:   Sample $\bm{\varepsilon}^{i}\sim\mathcal{N}(0,\sigma^{2}I)$
+17:   $\tilde{{\bm{x}}}^{i}\leftarrow\mathcal{P}({\bm{x}}+\bm{\varepsilon}^{i})$
+18:   $\tilde{{\bm{z}}}^{i}\leftarrow f_{e}(\tilde{{\bm{x}}}^{i})$
+19:   $\tilde{{\bm{z}}}^{i}_{m}\leftarrow\tilde{{\bm{z}}}^{i}+\mathcal{M}(\tilde{{\bm{z}}}^{i},\sigma)$
+20:  end for
+21:  Compute $l^{\mathcal{M}}_{\mathrm{rb}}$, $l_{\mathrm{id}}$, and $l_{\mathrm{stats}}$ using Eqs. 11 and 12.
+22:  Update $\mathcal{M}$ by gradient descent on $\nabla_{\mathcal{M}}\mathcal{L}_{\mathcal{M}}$
+23:  end for
+4.3 Further Discussion on Plugging GSB
+Certified robustness for the encoder $f_{e}$ with GSB. Let $f^{\prime}_{e}$ denote the feature encoder integrated with the proposed GSB. Then the forward process can be formulated as:
 $\vskip-2.84526pt{f}^{\prime}_{e}(\bm{x}+\bm{\varepsilon})=f_{e}(\mathcal{P}({\bm{x}}+\bm{\varepsilon}))+\mathcal{M}(\tilde{{\bm{z}}},\sigma).$ (14)
@@
 Where $\hat{S}^{\prime}(\bm{x})\in[0,1]$.
-We can prove that the Lipschitz property derived in Lemma 1 still holds, and the theoretical bound in Section 3 remains valid.
-Then, utilizing Theorem 1 and Corollary 1, we can derive the certified lower bound on $\mathrm{Cos}(\hat{f}^{\prime}_{e}(\bm{x^{\prime}}),f_{e}(\bm{x}))$ for any adversarial input $\bm{x}$$^{\prime}$.
-Algorithm 1 Training algorithm of PSM
-1:Input: dataset $\mathcal{D}$, feature encoder $f_{e}$, purifier $\mathcal{P}$, mapper $\mathcal{M}$, sampling number $n_{0}$, noise standard deviation $\sigma$, and loss weights $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$.
-2:Output: trained purifier $\mathcal{P}$ and mapper $\mathcal{M}$.
-3:// Stage 1: Fine-tune the purifier $\mathcal{P}$
-4:for each batch ${\bm{x}}\sim\mathcal{D}$ do
-5:  ${\bm{z}}=f_{e}({\bm{x}})$
-6:  for $i=1,\dots,n_{0}$ do
-7:    Sample $\bm{\varepsilon}^{i}\sim\mathcal{N}(0,\sigma^{2}I)$
-8:    $\tilde{{\bm{x}}}^{i}=\mathcal{P}({\bm{x}}+\bm{\varepsilon}^{i})$, and $\tilde{{\bm{z}}}^{i}=f_{e}(\tilde{{\bm{x}}}^{i})$
-9:  Compute $l^{\mathcal{P}}_{\mathrm{rb}}$ and $l_{\mathrm{mse}}$ using Equations 78.
-10:  Update $\mathcal{P}$ by gradient descent on $\nabla_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}$
-11:// Stage 2: Train the smoothness mapper $\mathcal{M}$
-12:for each batch ${\bm{x}}\sim\mathcal{D}$ do
-13:  ${\bm{z}}=f_{e}({\bm{x}})$
-14:  for $i=1,\dots,n_{0}$ do
-15:    Sample $\bm{\varepsilon}^{i}\sim\mathcal{N}(0,\sigma^{2}I)$
-16:    $\tilde{{\bm{x}}}^{i}=\mathcal{P}({\bm{x}}+\bm{\varepsilon}^{i})$, and $\tilde{{\bm{z}}}^{i}_{m}=f_{e}(\tilde{{\bm{x}}}^{i})+\mathcal{M}(\tilde{{\bm{z}}}^{i},\sigma)$
-17:  Compute $l^{\mathcal{M}}_{\mathrm{rb}}$, $l_{\mathrm{id}}$ and $l_{\mathrm{stats}}$ using Equations 1112.
-18:  Update $\mathcal{M}$ by gradient descent on $\nabla_{\mathcal{M}}\mathcal{L}_{\mathcal{M}}$
-Table 1: Experimental results on adversarial robustness of different defense methods and MLLMs on image captioning tasks.
-Values in parentheses denote the average textual similarity measured by GPTScore.
-The overall best results are shown in bold, and the best results without smoothing are underlined, highlighting the significant performance gain introduced by our proposed FS-PSM.
-Model Method M-Attack [23] FOA [14] AttackVLM [51]
-[…]
-LLaVA- 1.5-7B Org. 0.385 1% (0.06) 93% (0.54) 0.388 1% (0.049) 94% (0.578) 0.430 3% (0.089) 88% (0.490)
-Smoothed org. 0.650 82% (0.694) 2% (0.028) 0.652 87% (0.735) 1% (0.021) 0.660 89%(0.727) 2% (0.018)
-FARE [37] 0.588 44% (0.409) 24% (0.097) 0.504 19% (0.197) 51% (0.222) 0.499 18%(0.190) 44%(0.221)
-Smoothed FARE 0.716 59% (0.501) 14% (0.069) 0.656 35% (0.327) 29% (0.153) 0.729 60% (0.503) 11% (0.061)
-TeCoA [29] 0.720 51% (0.458) 16% (0.081) 0.674 21% (0.236) 37% (0.187) 0.606 17% (0.179) 43% (0.218)
-Smoothed TeCoA 0.805 59% (0.498) 14% (0.056) 0.789 39% (0.364) 20% (0.094) 0.755 42%(0.358) 18% (0.104)
-Open Flamingo -9B Org. 0.351 1%(0.089) 86%(0.583) 0.347 1%(0.091) 87%(0.569) 0.442 16%0.222) 59%(0.379)
-Smoothed org. 0.673 73%(0.655) 1%(0.063) 0.672 69%(0.623) 0%(0.061) 0.703 73%(0.687) 0%(0.055)
-FARE [37] 0.588 36%(0.407) 17%(0.145) 0.504 20%(0.249) 40%(0.240) 0.499 15%(0.243) 40%(0.247)
-Smoothed FARE 0.716 55%(0.529) 9%(0.088) 0.656 40%(0.414) 17%(0.125) 0.729 50%(0.488) 4%(0.089)
-TeCoA [29] 0.720 51%(0.472) 13%(0.116) 0.674 21%(0.240) 30%(0.210) 0.606 19%(0.244) 34%(0.225)
-Smoothed TeCoA 0.805 56%(0.518) 5%(0.090) 0.789 50%(0.454) 8%(0.095) 0.755 38%(0.398) 12%(0.126)
+We can prove that the Lipschitz property derived in Lemma 3.1 still holds, and the theoretical bound in Section 3 remains valid.
+Then, utilizing Theorem 3.2 and Corollary 3.3, we can derive the certified lower bound on $\mathrm{Cos}(\hat{f}^{\prime}_{e}(\bm{x^{\prime}}),f_{e}(\bm{x}))$ for any adversarial input $\bm{x}$$^{\prime}$.
 Training algorithm and dataset construction.
-The training procedure of PSM is summarized in Algorithm 1, where the purifier $\mathcal{P}$ and the smoothness mapper $\mathcal{M}$ are trained sequentially via a two-stage manner.
+The training procedure of GSB is summarized in Algorithm 1, where the denoiser $\mathcal{P}$ and the smoothness mapper $\mathcal{M}$ are trained sequentially via a two-stage manner.
 The expectation over Gaussian perturbations is approximated by Monte Carlo sampling with $n_{0}$ samples drawn from $\mathcal{N}(0,\sigma^{2}I)$, where we set $n_{0}=8$ in our training to balance efficiency and estimation accuracy.
-Notably, all modules can be trained in a self-supervised fashion, without relying on task-specific labels.
-Thus, to promote generalization across diverse visual domains and ensure compatibility with various MLLMs’ downstream tasks, we collect the training dataset $\mathcal{D}$ comprising 5,000 images from heterogeneous domains, including medical, cartoon, and natural images.
+5 Certified Robustness via FS
+5.1 Feature-wise Certification
 Table 1: The average FCSB at different predetermined adversarial bounds $\epsilon$ and the average certified radius $\mathcal{R}$ for $\mathrm{FCSB}\geq 0.5$.
@@
 0.125 0.25 0.375 0.50 0.75
-CLIP-B16 0.25 0.828 0.623 0.313 -0.06 / 0.31
+CLIP-L14 0.25 0.828 0.623 0.313 -0.06 / 0.31
+CLIP-L14+GSB 0.907 0.763 0.508 0.152 / 0.38
+CLIP-L14 0.50 0.717 0.586 0.439 0.253 -0.132 0.33
+CLIP-L14+GSB 0.924 0.873 0.799 0.698 0.409 0.69
+To evaluate the effectiveness of the proposed FS framework with GSB, we adopt CLIP-L14 (Radford et al., 2021) as the vanilla feature encoder $f_{e}$ and assess the feature-wise certified robustness under the FS.
+We report both the certified FCSB under different adversarial constraints $\epsilon$ and the average certified radius $\mathcal{R}$ for $\mathrm{FCSB}\geq 0.5$.
+The results are shown in Table 1.
+For implementation, we randomly sample 500 images from the ImageNet dataset (Deng et al., 2009) and approximate the Gaussian expectation using n=10,000n=10{,}000 Monte Carlo samples.
+As shown in Table 1, integrating GSB consistently enhances both the certified FCSB across a range of perturbation magnitudes and the average certified radius $\mathcal{R}$, indicating substantial improvements in certified robustness.
+5.2 Prediction-wise Certification
+Beyond the feature-wise certification, the guarantees provided by FS can be propagated to the prediction space by smoothing the deepest-layer features with a solvable classification head, which enables a closed-form translation from feature-space certification to prediction-level robustness.
+To this end, we adopt CLIP-L14 with GSB as the last-layer feature extractor for downstream classification.
+Within the FS framework, we employ a learnable prototype-based cosine classification head, where each class is represented by a unit-norm prototype, and predictions are determined by cosine similarity in the normalized feature space.
+This design admits a direct characterization of class decision boundaries, allowing the certified feature-wise cosine similarity guaranteed by FS to be converted into certified prediction bounds.
+We also compared the performance with Randomized Smoothing (RS) (Cohen et al., 2019), where we utilize the same CLIP-L14+GSB as the feature encoder and train a learnable linear classification head for a fair comparison.
+In the experiments, we randomly sample 500 training instances from 10 ImageNet classes (Deng et al., 2009) to train both classification heads and evaluate them on an additional 500 test samples.
+For both FS and RS, the Gaussian expectation is approximated using n=10,000n=10{,}000 Monte Carlo samples.
+The comparative results in Table 2 demonstrate the effectiveness of the FS in providing certified prediction-level robustness.
+Further illustration and proof of prediction-wise certification via FS are provided in Appendix A.3.
+Table 2: The comparison results of the RS and FS framework on prediction-wise certification.
+Smoothing $\sigma$ Certified accuracy at different $\ell_{2}$ adv-bound $\varepsilon$ (%)
+0.00 0.125 0.25 0.375 0.5 0.75 1 1.50
+RS $0.25$ 93.4 88 86.2 81.0 75.4 66.5 / /
+FS 94.2 92.4 90.0 88.0 85.8 72.0 / /
+RS $0.50$ 91.2 86 84.4 81.4 77.0 71.6 67.0 49.2
+FS 91 85.2 84.0 83.6 83.0 77.6 69.2 54.0
+6 Experimental Results on MLLM Protection
+While end-to-end robustness certification for MLLMs is appealing and practically meaningful, smoothing the entire model and certifying its autoregressive predictions with a solvable head are both computationally expensive and challenging.
+Alternatively, we thereby seek the feature-wise certification of the visual encoder (e.g., CLIP) in MLLMs, and empirically evaluate the resulting task-level robustness under strong white-box adversarial attacks, providing a practical compromise between certification feasibility and effectiveness.
+Notably, we find that this feature-wise certification offers two advantages: (i) efficiency, as smoothing the lightweight encoder $f_{e}$ incurs substantially lower computational overhead than smoothing the full model $\mathcal{F}$; and (ii) generality, where feature-representation level certification makes it applicable to various downstream tasks.
+Table 3: Experimental results on adversarial robustness of different defense methods and MLLMs on image captioning tasks.
+Values in parentheses denote the average textual similarity measured by GPTScore.
+The overall best results are shown in bold, and the best results without smoothing are underlined, highlighting the significant performance gain introduced by our proposed FS.
+Model Method M-Attack (Li et al., 2025b) FOA (Jia et al., 2025) AttackVLM (Zhao et al., 2023)
 […]
-CLIP-B16 0.50 0.717 0.586 0.439 0.253 -0.132 0.33
-[…]
-Certified robustness bound of FS-PSM.
-To evaluate the effectiveness of the proposed PSM module, we adopt CLIP-B16 [33] as the vanilla feature encoder $f_{e}$ and assess the certified robustness bound of both the original smoothed encoder $\hat{f}_{e}$ and its PSM-enhanced counterpart $\hat{f}^{\prime}_{e}$ under the FS.
-We report both the certified FCSB under different adversarial constraints $\epsilon$ and the certified radius $\mathcal{R}$ for $\mathrm{FCSB}\geq 0.5$.
-The results are shown in Table 2.
-For implementation, we randomly sample 100 images from the ImageNet dataset [6] and approximate the Gaussian expectation using n=1,000n=1{,}000 Monte Carlo samples.
-As shown in Table 2, integrating PSM consistently enhances both the certified FCSB across a range of perturbation magnitudes and the average certified radius $\mathcal{R}$, indicating substantial improvements in certified robustness.
-To further assess the practical protection of our FS-PSM, we then conduct extensive experiments across multiple MLLMs and downstream tasks in the next section, evaluating their performance before and after smoothing under various strong white-box attacks.
-5 Experiments
+LLaVA- 1.5-7B Org. 0.385 1% (0.06) 93% (0.54) 0.388 1% (0.049) 94% (0.578) 0.430 3% (0.089) 88% (0.490)
+Smoothed org. 0.588 63% (0.515) 3% (0.027) 0.587 65% (0.519) 6% (0.028) 0.598 71%(0.556) 4% (0.03)
+FARE (Schlarmann et al., 2024) 0.588 44% (0.409) 24% (0.097) 0.504 19% (0.197) 51% (0.222) 0.499 18%(0.190) 44%(0.221)
+Smoothed FARE 0.687 64% (0.541) 10% (0.053) 0.634 43% (0.391) 14% (0.065) 0.653 52% (0.443) 7% (0.045)
+TeCoA (Mao et al., 2023) 0.720 51% (0.458) 16% (0.081) 0.674 21% (0.236) 37% (0.187) 0.606 17% (0.179) 43% (0.218)
+Smoothed TeCoA 0.801 61% (0.510) 17% (0.081) 0.769 39% (0.364) 10% (0.0530) 0.720 38% (0.355) 14% (0.065)
+Open Flamingo -9B Org. 0.351 1% (0.089) 86% (0.583) 0.347 1% (0.091) 87%(0.569) 0.442 16% (0.222) 59% (0.379)
+Smoothed org. 0.592 59% (0.483) 3% (0.081) 0.588 57% (0.472) 1% (0.066) 0.703 61% (0.505) 0% (0.051)
+FARE (Schlarmann et al., 2024) 0.588 36% (0.407) 17% (0.145) 0.504 20% (0.249) 40% (0.240) 0.499 15% (0.243) 40% (0.247)
+Smoothed FARE 0.782 48% (0.462) 5% (0.062) 0.768 35% (0.354) 9% (0.102) 0.754 40% (0.404) 5% (0.086)
+TeCoA (Mao et al., 2023) 0.720 51% (0.472) 13% (0.116) 0.674 21% (0.240) 30% (0.210) 0.606 19% (0.244) 34% (0.225)
+Smoothed TeCoA 0.835 48% (0.464) 3% (0.067) 0.818 39% (0.371) 5% (0.088) 0.805 35%(0.341) 8%(0.079)
 6.1 Experimental Setup
 Evaluated models and tasks.
-Since the proposed FS requires access to the forward feature computation process of MLLMs, we primarily validate its effectiveness on open-sourced MLLMs, including LLaVA-V1.5-7B [24, 25] and OpenFlamingo9B [1].
-We comprehensively assess the performance of plugging the FS-PSM under adversarial conditions across multiple downstream tasks, including:
-  • Image captioning: Following the setup in [23, 14], we randomly take 100 images from the NIPS 2017 Adversarial Attacks and Defenses Competition dataset 5.1 and ask the model to caption the image.
-  • Image Classification: We randomly sample 500 images from 10 classes in the ImageNet dataset [6] and ask the model to classify the input.
-  • Visual Question Answering (VQA): We utilize 100 image-question pairs from the ScienceQA dataset [26].
+Since the proposed FS requires access to the forward feature computation process of MLLMs, we primarily validate its effectiveness on open-sourced MLLMs, including LLaVA-V1.5-7B (Liu et al., 2023, 2024) and OpenFlamingo9B (Awadalla et al., 2023).
+We comprehensively assess the performance of plugging the FS under adversarial conditions across multiple downstream tasks, including:
+  • Image captioning: Following (Li et al., 2025b; Jia et al., 2025), we randomly take 100 images from the NIPS 2017 Adversarial Attacks and Defenses Competition dataset 6.1 and ask the model to caption the image.
+  • Image Classification: We randomly sample 500 images from 10 classes in the ImageNet dataset (Deng et al., 2009) and ask the model to classify the input.
+  • Visual Question Answering (VQA): We utilize 100 image-question pairs from the ScienceQA dataset (Lu et al., 2022) and ask the model to select the answer.
 footnotetext: https://nips.cc/Conferences/2017/CompetitionTrack
-Meanwhile, we also evaluate our FS-PSM on the original CLIP-L14 [33] on image classification for the generality.
-Table 3: Experimental results on adversarial robustness of different defense methods and MLLMs on image classification tasks.
-The overall best results are shown in bold, and the best results without smoothing are underlined.
-Model Method M-Attack [23] FOA [14] AttackVLM [51]
-[…]
-LLaVA- 1.5-7B Org. 0.427 8.2% 78.2% 0.437 3.8 % 81.0 % 0.458 6.0% 78.4%
-Smoothed org. 0.681 90.4% 0.6% 0.697 88.8% 0.4% 0.700 88.4% 0.2%
-FARE [37] 0.574 55.4% 14.6% 0.521 24.8% 39.8% 0.508 25.8% 45.0%
-Smoothed FARE 0.859 75.6% 5.2% 0.712 72.0% 5.6% 0.699 74.0% 6.2%
-TeCoA [29] 0.731 67.4% 2% 0.626 29.0% 20.0% 0.592 33.6% 17.8%
-Smoothed TeCoA 0.819 74.4% 0.4% 0.727 57.2% 3.0% 0.711 57.4% 2.0%
-CLIP -L14 Org. 0.427 18.0% 37.2% 0.437 10.0 % 52.4 % 0.458 16.0% 48.4%
-Smoothed org. 0.681 92.4% 0.4% 0.697 92.0% 0.4% 0.700 92.0% 0.4%
-FARE [37] 0.574 70.4% 5.2% 0.521 42.0% 14.4% 0.508 35.6% 24.4%
-Smoothed FARE 0.859 85.2% 0.8% 0.712 81.6% 0% 0.699 82.4% 0%
-TeCoA [29] 0.731 81.2% 2.4% 0.626 39.2% 20.2% 0.592 35.8% 24.4%
-Smoothed TeCoA 0.819 88.4% 1.6% 0.727 76.0% 4.8% 0.711 73.2% 6.0%
 The threat model.
-To comprehensively assess the robustness under strong adversaries, we employ three SOTA adversarial attacks specifically designed for MLLMs, named AttackVLM [51], M-Attack [23], and FOA [14] using the white-box setting.
+To comprehensively assess the robustness under strong adversaries, we employ three SOTA adversarial attacks specifically designed for MLLMs, named AttackVLM (Zhao et al., 2023), M-Attack (Li et al., 2025b), and FOA (Jia et al., 2025) using the white-box setting, where attackers can fully access the feature extractor and GSB.
 All attacks are implemented following their original best configurations, with the adversarial perturbation budget $\epsilon$ set to $\|\epsilon\|_{\infty}=16/255$.
-We also consider a stronger bound with $\|\epsilon\|_{\infty}=32/255$ for a stress test.
 Compared defense methods.
-As this work pioneers the research for MLLMs’ certified robustness, we primarily compare our defense with adversarial training based approaches, namely FARE [37] and TeCoA [29].
+As robustness certification for MLLMs remains largely unexplored, we primarily compare our method against adversarial training–based defenses, specifically FARE (Schlarmann et al., 2024) and TeCoA (Mao et al., 2023).
 Both FARE and TeCoA adopt adversarial training to obtain robust feature encoders that can be directly integrated into models such as LLaVA-1.5-7B and OpenFlamingo-9B.
@@
 Implementation details.
-For practical inference efficiency, we set the number of samples to $n_{0}=4$ for smoothing, which provides a favorable trade-off between robustness and runtime.
-We train three independent PSM modules, utilizing the vanilla feature encoders from models including LLaVA-1.5-7B, OpenFlamingo-9B, and CLIP-L14.
-To further assess cross-model generalization, we evaluate the PSM trained on a vanilla encoder by directly applying it to adversarially trained encoders FARE and TeCoA.
-For all feature encoders, we set hyperparameters $\lambda_{1},~\lambda_{2},~\lambda_{4}=0.25$, $\lambda_{3}=100$, and $\sigma=0.25$.
-In our tables, the term ”smoothed” denotes the process of smoothing the encoder via FS and enhancing it with PSM.
+For practical inference efficiency, we set the number of samples to $n_{0}=8$ for smoothing, ensuring a favorable trade-off between robustness and runtime.
+We train three independent GSB modules for different models, including the encoder of LLaVA-1.5-7B, the encoder of OpenFlamingo-9B, and CLIP-L14.
+To further assess cross-model generalization, we directly utilize the GSB trained on a vanilla encoder on adversarially trained encoders, including FARE and TeCoA.
+For all feature encoders, we set hyperparameters $\lambda_{1},~\lambda_{3}=0.25$, $\lambda_{2}=100$, and $\sigma=0.25$.
+In our tables, the term ”smoothed” denotes the process of smoothing the encoder via FS and enhancing it with GSB.
 Evaluation metrics.
 We mainly report: the Feature Cosine Similarity (FCS), the Accuracy (ACC), and the Attack Success Rate (ASR).
-Specifically, FCS measures the cosine similarity between clean and adversarial features extracted by the same feature encoder, reflecting its feature-space robustness.
+Specifically, FCS measures the cosine similarity between clean and adversarial features extracted by the same feature encoder, reflecting its feature-wise robustness.
 ACC denotes the proportion of correctly completed tasks, while ASR indicates the proportion of cases where the model is successfully manipulated to produce the adversarially targeted outputs.
 For image classification and VQA tasks, ACC and ASR are determined by whether the MLLM outputs match the correct or adversarial targets.
-For image captioning, following [23], we adopt the LLM-as-a-judge protocol [52] to evaluate ACC and ASR.
+For image captioning, following (Li et al., 2025b), we adopt the LLM-as-a-judge protocol (Zheng et al., 2023) to evaluate ACC and ASR.
 Specifically, we first generate the clean and adversarially targeted captions by feeding the clean and targeted inputs into the vanilla MLLM (e.g., the original LLaVA).
 We then obtain the predicted caption by feeding the adversarially perturbed input into the tested model (e.g., Smoothed LLaVA).
-The textual similarity is computed using GPTScore [9], where a task is considered successful if the GPTScore between the predicted and clean captions is $\mathbf{\geq 0.5}$, and an attack is considered successful if the GPTScore between the predicted and adversarially targeted captions is $\mathbf{\geq 0.3}$.
+The textual similarity is computed using GPTScore (Fu et al., 2024), where a task is considered successful if the GPTScore between the predicted and clean captions is $\mathbf{\geq 0.5}$, and an attack is considered successful if the GPTScore between the predicted and adversarially targeted captions is $\mathbf{\geq 0.3}$.
+Table 4: Experimental results on adversarial robustness of different defense methods and MLLMs on image classification tasks.
+The overall best results are shown in bold, and the best results without smoothing are underlined.
+Model Method M-Attack (Li et al., 2025b) FOA (Jia et al., 2025) AttackVLM (Zhao et al., 2023)
+[…]
+LLaVA- 1.5-7B Org. 0.427 8.2% 78.2% 0.437 3.8 % 81.0 % 0.458 6.0% 78.4%
+Smoothed org. 0.590 84.8% 0.2% 0.600 87.2% 0.2% 0.605 87.0% 0.2%
+FARE (Schlarmann et al., 2024) 0.574 55.4% 14.6% 0.521 24.8% 39.8% 0.508 25.8% 45.0%
+Smoothed FARE 0.695 75.6% 5.2% 0.681 66.2% 1% 0.671 68.0% 1%
+TeCoA (Mao et al., 2023) 0.731 67.4% 2% 0.626 29.0% 20.0% 0.592 33.6% 17.8%
+Smoothed TeCoA 0.790 70.6% 0.6% 0.714 58.8% 0.8% 0.712 56.4% 0.4%
 Table 5: Experimental results on adversarial robustness of different defense methods and MLLMs on VQA tasks.
 The overall best results are shown in bold, and the best results without smoothing are underlined.
-Model Method M-Attack [23] FOA [14] AttackVLM [51]
+Model Method M-Attack (Li et al., 2025b) FOA (Jia et al., 2025) AttackVLM (Zhao et al., 2023)
 […]
 LLaVA- 1.5-7B Org. 0.398 31% 28% 0.383 22% 22% 0.474 25% 27%
-Smoothed org. 0.732 47% 0% 0.676 43% 0% 0.765 43% 1%
-FARE [37] 0.657 47% 5% 0.590 32% 0% 0.550 38% 7%
-Smoothed FARE 0.859 48% 1% 0.768 38% 0% 0.818 39% 0%
-TeCoA [29] 0.788 31% 2% 0.748 31% 2% 0.667 31% 1%
-Smoothed TeCoA 0.917 32% 1% 0.858 34% 1% 0.862 29% 1%
+Smoothed org. 0.682 47% 0% 0.643 43% 0% 0.712 43% 1%
+FARE (Schlarmann et al., 2024) 0.657 47% 5% 0.590 32% 0% 0.550 38% 7%
+Smoothed FARE 0.861 48% 1% 0.788 38% 0% 0.825 39% 0%
+TeCoA (Mao et al., 2023) 0.788 31% 2% 0.748 31% 2% 0.667 31% 1%
+Smoothed TeCoA 0.925 32% 1% 0.880 34% 1% 0.879 29% 1%
 6.2 Experimental Results on Different Tasks
 Image captioning.
-The comparative results of different defense methods and MLLMs on the image captioning task are presented in Table 1.
-In this task, the adversaries aim to manipulate the MLLMs into producing the caption corresponding to a completely unrelated and maliciously chosen target image.
-These results demonstrate that converting MLLMs into their smoothed variants via the proposed FS-PSM yields consistently strong robustness across diverse adversarial attacks, while empirical defenses exhibit a noticeable performance drop under stronger attacks.
-When adversaries vary from M-Attack to FOA, the accuracy of LLaVA with FARE drops from 44% to 19%, and that of TeCoA declines from 51% to 21%.
-Notably, utilizing FS-PSM achieves remarkable improvements, with the accuracy of LLaVA increasing from 1% to 87% and the ASR dropping from 94% to 1% under the strongest FOA attack.
-Furthermore, the results demonstrate that FS-PSM possesses strong cross-model generalization; directly integrating it into FARE and TeCoA without any fine-tuning consistently enhances their robustness.
+Table 3 reports the image captioning results of MLLMs under different attacks and defenses.
+The adversaries aim to manipulate the MLLMs into producing the caption corresponding to a maliciously chosen target image.
+The results show that converting MLLMs into smoothed variants via the proposed FS smoothing yields consistently strong robustness across diverse attacks, whereas empirical defenses degrade substantially under stronger adversaries.
+When attacks escalate from M-Attack to FOA, the accuracy of LLaVA with FARE drops from 44% to 19%, and that of TeCoA from 51% to 21%.
+In contrast, FS achieves significant robustness gains, increasing LLaVA accuracy from 1% to 65% and reducing the ASR from 94% to 6% under the strongest FOA attack.
+Moreover, FS exhibits strong cross-model generalization, consistently improving the robustness of FARE and TeCoA without additional fine-tuning.
 Image classification.
-The comparative results of different defense methods on the image classification are presented in Table 3.
-The adversarial objective is to mislead the model into classifying an adversarial image into a maliciously targeted class that is semantically unrelated to the original one.
-For image classification with CLIP-L14, we train a one-layer classification head on top of its extracted features.
-The results indicate that FS-PSM provides consistent and substantial robustness improvements for classification tasks.
-Meanwhile, those adversarial training-based defenses exhibit a noticeable performance drop under strong FOA attacks.
-The accuracy of FARE with LLaVA drops from 55.4% to 24.8%, and that of TeCoA declines from 67.4% to 29.0% when adversaries vary from M-Attack to FOA.
-Integrating FS-PSM into the vanilla model achieves remarkable improvements, with the accuracy of LLaVA increasing from 3.8% to 88.0% and the ASR dropping from 81% to 0.4% under the strongest FOA attack.
-Meanwhile, directly integrating the FS-PSM brings huge robustness gains for those adversarially trained feature encoders.
+Table 4 presents the classification results, indicating that FS delivers consistent and substantial robustness improvements across all settings.
+The adversarial objective is to mislead the model into classifying an adversarial image into a maliciously targeted class.
+integrating FS into the vanilla model increases LLaVA accuracy from 3.8% to 87.2% and reduces the ASR from 81% to 0.2% under the strongest FOA attack.
+Similar robustness gains are observed when FS is applied to FARE and TecoA.
 VQA.
 The comparative results of different defense methods on LLaVA-1.5-7B on the VQA are presented in Table 5.
-The performances of OpenFlamingo and CLIP are omitted, as these models are not directly applicable to the VQA task.
 In this setting, the adversarial objective is to mislead the model into selecting a wrong option, ”None of the above”, for each image–question pair.
-This task presents a greater challenge for pure vision-based adversaries than the previous two, as MLLMs can often infer correct answers directly from textual cues without relying heavily on visual inputs.
-The results demonstrate that incorporating our method yields substantial performance improvements across all attack types, significantly enhancing prediction accuracy while reducing the attack success rate to nearly zero.
-More implementation details for the main results are presented in our supplementary material, section S.3.
-5.3 Further analysis on the main results
-Reasons for FARE and TeCoA performing poorly with high FCS.
-Although those encoders preserve high FCS with respect to their clean inputs, the adversarial training process inherently induces a distributional shift in the feature representations.
-Subsequent adversarial perturbations further amplify this shift, leading to more pronounced performance degradation.
-In addition, the limited diversity of adversarial training data restricts their feature generalization ability, amplifying the performance drop when evaluated on downstream tasks with unseen or mismatched data distributions.
-Advantages of our proposed FS with PSM.
-The experimental results highlight three key advantages: (1) Effectiveness: utilizing the proposed FS-PSM substantially enhances the performance of various MLLMs under adversarial attacks; (2) Trustworthiness: compared with adversarial training, FS-PSM provides more stable and consistently stronger protection across a wide range of attacks; (3) Generality: FS-PSM can be combined with other defenses to enhance the effectiveness without additional fine-tuning.
-5.4 Ablation study
-Table 5: The ablation study.
-We report the average FCSB and the average $\mathcal{R}$ for $\mathrm{FCSB}\geq 0.5$ for certified robustness.
-We report the average accuracy and ASR for empirical robustness under FOA attack.
-The clean accuracy without attack is 92.4%.
-Encoder $\sigma$ Avg. FCSB at different $\epsilon$ Avg. $\mathcal{R}$ Acc ASR
+This task presents a greater challenge for pure vision-based adversaries than the previous two, as MLLMs can often infer correct answers directly from textual cues.
+The results demonstrate that incorporating FS consistently yields substantial performance improvements across all attack types, significantly improving prediction accuracy while driving the ASR to nearly zero.
+Reasons for the poor performance of FARE and TeCoA despite high FCS.
+Although these encoders preserve high FCS relative to their clean inputs, adversarial training inherently induces a shift in the feature-space distribution.
+Subsequent adversarial perturbations and the limited diversity of adversarial training data further exacerbate this mismatch, leading to pronounced performance degradation on downstream tasks with unseen data distributions.
+6.3 Ablation Study and Analysis of Efficiency
+Table 6: The ablation study. The clean accuracy is 91.0%.
+Smoothing Encoder $\sigma$ Avg. FCSB at different $\epsilon$ Avg. $\mathcal{R}$ Acc
 0.25 0.50 0.75
-CLIP-B16 / / / / / 1.6% 95.6%
-FS CLIP-B16 0.50 0.586 0.253 -0.132 0.33 42.4% 0.8%
+/ CLIP-L14 / / / / / 16.0%
+FS CLIP-L14 0.50 0.662 0.360 -0.018 0.41 66.2%
 […]
 […]
-To rigorously assess the contribution of the proposed FS and each module of PSM, we adopt CLIP-B16 as the base feature encoder and evaluate the performance on the image classification task.
-We report both the theoretically certified robustness and the empirical performance under the strong FOA attack.
-The results are summarized in Table 5, where the first row indicates the performance of the vanilla CLIP-B16 model without utilizing the FS.
-The next three rows are the results of the smoothed CLIP-B16 with FS, smoothed CLIP-B16 with FS and mapper $\mathcal{M}$, and smoothed CLIP-B16 with FS and PSM.
-The results demonstrate that enforcing smoothness in the feature space substantially enhances practical robustness against adversarial perturbations, raising the accuracy from 1.6% to 91.6% and reducing the ASR from 95.6% to 0.4%.
-Meanwhile, each component makes a significant contribution to improving both the certified radius $\mathcal{R}$ and adversarial robustness.
-More experimental results on: 1) the performance of using different purifier (a lightweight U-Net denoiser), 2) the performance under a stronger adversarial bound $\|\epsilon\|_{\infty}=32$, and 3) efficiency–robustness trade-off under various $n_{0}$, are provided in our supplementary material, section S.3.
+[…]
+Ablation: To rigorously assess the contribution of the proposed FS and each module of GSB (including denoiser $\mathcal{P}$ and mapper $\mathcal{M}$), we adopt CLIP-L14 as the base feature encoder and conduct the image classification task following the design in Section 5.2.
+We report both the feature-wise certification and the empirical performance under the FOA attack.
+The results are summarized in Table 6, where the first row indicates the performance of the vanilla CLIP-L14 without FS.
+The next four rows are the results of the smoothed CLIP-L14 with different modules enhanced.
+The results demonstrate that each component makes a significant contribution to improving both the certified radius $\mathcal{R}$ and adversarial robustness.
+Meanwhile, it demonstrates that enforcing smoothness in the feature space substantially enhances practical robustness against adversarial perturbations, raising the accuracy from 16.0% to 89.6%.
+Table 7: The analysis of efficiency.
+Model $n_{0}$ Avg. infer time (s) FCS
+LLaVA-1.5 7B 1 0.41 0.450
+4 0.52 0.562
+8 0.69 0.590
+64 3.08 0.609
+Efficiency analysis: Table 7 presents the efficiency analysis under varying numbers of Gaussian samples, evaluated on a single RTX 4090 GPU.
+We consider the image captioning task and report the average per-image inference time together with the average FCS under FOA attack.
+Compared to whole-model smoothing, feature-wise smoothing incurs lower inference overhead; notably, setting $n_{0}=8$ increases the per-image inference time of LLaVA by only $\sim$0.28 seconds, highlighting the practicality of FS for real-world deployment.
 7 Conclusion.
-This work pioneers the research on establishing the certified robustness of MLLMs via a feature-space perspective.
-By introducing the Feature-space Smoothing (FS) framework, we show how to transform any feature encoder into a smoothed version that is equipped with a theoretical lower bound on the cosine similarity between clean and adversarial representations.
-Moreover, we introduce the Purifier and Smoothness Mapper (PSM), a plug-and-play module that can be seamlessly integrated into existing MLLMs to amplify the certified robustness under FS without any fine-tuning.
-Extensive experiments demonstrate that the proposed FS-PSM not only provides strong theoretical robustness guarantees but also achieves superior empirical protection against various attacks.
-Supplementary Material
-S.7 Proof of core theorem
-S.7.1 Proof of lemma1
+This work pioneers the research on establishing the feature-space certified robustness of MLLMs.
+By introducing the Feature-space Smoothing (FS) framework, we show how to transform a given encoder into a smoothed version that is equipped with a theoretical lower bound on the cosine similarity between clean and adversarial representations.
+Moreover, we propose the Gaussian Robustness Booster (GSB), a plug-and-play module that seamlessly integrates into existing MLLMs to enhance the certified robustness guaranteed by FS.
+Extensive experiments demonstrate that FS provides effective adversarial protection for MLLMs in a plug-and-play manner while incurring modest inference latency, making it promising for real-world deployment on MLLM protection.
+Appendix
+Appendix A.1 Proof of core theorem
+A.1.1 Proof of lemma1
 Note that:
-$\nabla\Phi^{-1}(\hat{S}_{x_{c}}(\bm{x}))=\frac{\nabla\hat{S}_{x_{c}}(\bm{x})}{\Phi^{\prime}\bigr(\Phi^{-1}(\hat{S}_{x_{c}}(\bm{x}))\bigr)}.$ (S.16)
+$\nabla\Phi^{-1}(\hat{S}_{x_{c}}(\bm{x}))=\frac{\nabla\hat{S}_{x_{c}}(\bm{x})}{\Phi^{\prime}\bigr(\Phi^{-1}(\hat{S}_{x_{c}}(\bm{x}))\bigr)}.$ (A.16)
 Denote $p=\hat{S}_{{\bm{x}}_{c}}({\bm{x}})$, we can get that:
-$\Phi^{\prime}\bigr(\Phi^{-1}(\hat{S}_{x_{c}}(\bm{x}))\bigr)=\frac{1}{\sqrt{2\pi}}\exp\!\bigl(-\tfrac{1}{2}(\Phi^{-1}(p))^{2}\bigr)$ (S.17)
+$\Phi^{\prime}\bigr(\Phi^{-1}(\hat{S}_{x_{c}}(\bm{x}))\bigr)=\frac{1}{\sqrt{2\pi}}\exp\!\bigl(-\tfrac{1}{2}(\Phi^{-1}(p))^{2}\bigr)$ (A.17)
 Thus, we need to prove that for any unit direction $u$.
-$u\cdot\nabla\hat{S}_{{\bm{x}}_{c}}(\bm{x})\leq\frac{1}{\sqrt{2\pi}}\exp\!\bigl(-\tfrac{1}{2}(\Phi^{-1}(p))^{2}\bigr).$ (S.18)
+$u\cdot\nabla\hat{S}_{{\bm{x}}_{c}}(\bm{x})\leq\frac{1}{\sqrt{2\pi}}\exp\!\bigl(-\tfrac{1}{2}(\Phi^{-1}(p))^{2}\bigr).$ (A.18)
 For the left-hand side, we can get:
-$\nabla\hat{S}_{x_{c}}(\bm{x})=\frac{1}{(2\pi)^{n/2}}\int_{\mathbb{R}^{n}}S_{{\bm{x}}_{c}}(\bm{t})\,({\bm{x}}-{\bm{t}})\,\exp\!\bigl(-\tfrac{1}{2}\|{\bm{x}}-{\bm{t}}\|^{2}\bigr)\,\mathrm{d}{\bm{t}},$ (S.19)
+$\nabla\hat{S}_{x_{c}}(\bm{x})=\frac{1}{(2\pi)^{n/2}}\int_{\mathbb{R}^{n}}S_{{\bm{x}}_{c}}(\bm{t})\,({\bm{x}}-{\bm{t}})\,\exp\!\bigl(-\tfrac{1}{2}\|{\bm{x}}-{\bm{t}}\|^{2}\bigr)\,\mathrm{d}{\bm{t}},$ (A.19)
 which can also be rewritten as follows:
@@
 We can now claim that the supremum of the above quantity over all encoders $f:{\bm{x}}\to{\bm{z}}$, subject to the constraint that $p=\mathbb{E}[{S}_{x_{c}}(x+X)]$, is equal to:
-$\mathbb{E}\bigl[(X\cdot u)\,1\{X\cdot u\geq-\Phi^{-1}(p)\}\bigr]=\frac{1}{\sqrt{2\pi}}\exp\!\Bigl(-\tfrac{1}{2}\bigl(\Phi^{-1}(p)\bigr)^{2}\Bigr),$ (S.20)
+$\mathbb{E}\bigl[(X\cdot u)\,1\{X\cdot u\geq-\Phi^{-1}(p)\}\bigr]=\frac{1}{\sqrt{2\pi}}\exp\!\Bigl(-\tfrac{1}{2}\bigl(\Phi^{-1}(p)\bigr)^{2}\Bigr),$ (A.20)
 which concludes the proof of Lemma 1.
-S.7.2 Proof of Theorem1
+A.1.2 Proof of Theorem1
 From Lemma 1, we can get:
-$\left\|{\bm{x}}-{\bm{x}}^{\prime}\right\|\geq\Phi^{-1}(\hat{S}_{{\bm{x}}}({\bm{x}}))-\Phi^{-1}(\hat{S}_{{\bm{x}}}({\bm{x}}^{\prime})).$ (S.21)
+$\left\|{\bm{x}}-{\bm{x}}^{\prime}\right\|\geq\Phi^{-1}(\hat{S}_{{\bm{x}}}({\bm{x}}))-\Phi^{-1}(\hat{S}_{{\bm{x}}}({\bm{x}}^{\prime})).$ (A.21)
 As $\left\|{\bm{x}}-{\bm{x}}^{\prime}\right\|\leq\epsilon$, we can get:
-$\Phi^{-1}(\hat{S}_{{\bm{x}}}({\bm{x}}^{\prime}))\geq\Phi^{-1}(\hat{S}_{{\bm{x}}}({\bm{x}}))-\epsilon.$ (S.22)
+$\Phi^{-1}(\hat{S}_{{\bm{x}}}({\bm{x}}^{\prime}))\geq\Phi^{-1}(\hat{S}_{{\bm{x}}}({\bm{x}}))-\epsilon.$ (A.22)
 Meanwhile, as $f_{e}({\bm{x}})$ is a unit vector, using Equation 5 in the main paper, we can get that:
 $\hat{S}({\bm{x}})$ $=\tfrac{1}{2}\Big(1+\big\langle\mathbb{E}[f_{e}({\bm{x}}+\varepsilon)],\,f_{e}({\bm{x}})\big\rangle\Big)$
-$=\tfrac{1}{2}\Big(1+\langle\hat{f}_{e}({\bm{x}}),\,f_{e}({\bm{x}})\rangle\Big).$ (S.23)
+$=\tfrac{1}{2}\Big(1+\langle\hat{f}_{e}({\bm{x}}),\,f_{e}({\bm{x}})\rangle\Big).$ (A.23)
 where $\langle\cdot,\cdot\rangle$ is the inner product. Hence, we can get:
-$\langle\hat{f}_{e}({\bm{x}}),\,f_{e}({\bm{x}})\rangle=2\hat{S}({\bm{x}})-1.$ (S.24)
+$\langle\hat{f}_{e}({\bm{x}}),\,f_{e}({\bm{x}})\rangle=2\hat{S}({\bm{x}})-1.$ (A.24)
 As $\left\|\hat{f}_{e}({\bm{x}}^{\prime})\right\|\leq 1$ and $\left\|f_{e}({\bm{x}})\right\|=1$, we can get:
-${Cos}\big(\hat{f}({\bm{z}}),\,f_{e}({\bm{x}})\big)$ $=\frac{\langle\hat{f}({\bm{z}}),f_{e}({\bm{x}})\rangle}{\|\hat{f}({\bm{z}})\|}\;$ (S.25)
+${Cos}\big(\hat{f}({\bm{z}}),\,f_{e}({\bm{x}})\big)$ $=\frac{\langle\hat{f}({\bm{z}}),f_{e}({\bm{x}})\rangle}{\|\hat{f}({\bm{z}})\|}\;$ (A.25)
 $\geq\;\langle\hat{f}({\bm{z}}),f_{e}({\bm{x}})\rangle$
 By the monotonicity of $\Phi$,
-$\hat{S}_{{\bm{x}}}({\bm{x}}^{\prime})\;\geq\;\Phi\!\Big(\Phi^{-1}\!\big(\hat{S}({\bm{x}})\big)-\epsilon\Big).$ (S.26)
-Combining Equations S.24S.25S.26, we thereby can derive that:
-$\mathrm{Cos}\big(\hat{f}({\bm{x}}^{\prime}),\,f_{e}({\bm{x}})\big)\;\geq\;2\,\Phi\!\Big(\Phi^{-1}\!\big(\hat{S}({\bm{x}})\big)-\epsilon\Big)-1,$ (S.27)
+$\hat{S}_{{\bm{x}}}({\bm{x}}^{\prime})\;\geq\;\Phi\!\Big(\Phi^{-1}\!\big(\hat{S}({\bm{x}})\big)-\epsilon\Big).$ (A.26)
+Combining Equations A.24A.25A.26, we thereby can derive that:
+$\mathrm{Cos}\big(\hat{f}({\bm{x}}^{\prime}),\,f_{e}({\bm{x}})\big)\;\geq\;2\,\Phi\!\Big(\Phi^{-1}\!\big(\hat{S}({\bm{x}})\big)-\epsilon\Big)-1,$ (A.27)
 which concludes the proof. For noise variance $\sigma^{2}$, the same proof gives the bound with $\epsilon$ replaced by $\epsilon/\sigma$.
-S.8 Training details of PSM
-S.8.1 Details of fine-tuning noise purifier
-The Purifier $\mathcal{P}$ is optimized in a self-supervised manner using the objective $\mathcal{L}_{p}$ defined in Equation 9.
-We employ the ImageNet-pretrained $256\times 256$ unconditional diffusion modelS.8.1 as the backbone of $\mathcal{P}$.
-All input images are resized to $256\times 256$.
-To perform denoising, we instantiate a diffusion process consisting of 1,000 timesteps with a linear noise scheduler.
-Following common practices, the rescaled learned noise variance (i.e., ”rescale_learned_sigma”) is disabled to prevent instability during fine-tuning.
-We fine-tune the model for a total of 12,000 optimization steps using the AdamW optimizer, with an initial learning rate of $5\times 10^{-5}$.
-An exponential moving average (EMA) with a decay rate of 0.99 is maintained throughout training to stabilize updates and improve the effective generalization of the learned purifier.
-footnotetext: https://github.com/openai/guided-diffusion?tab=readme-ov-file
-Figure S.3: Structure of the residual smoothness mapper.
-S.8.2 Details of training residual mapper
-The structure of the proposed residual smoothness mapper $\mathcal{M}$ is illustrated in Figure S.3.
+Appendix A.2 Training details of GSB
+A.2.1 Details of Denoiser
+The denoiser $\mathcal{P}$ is implemented as a lightweight conditional U-Net and trained in a supervised denoising paradigm using a mean squared error (MSE) objective to regress clean images from their noisy counterparts.
+$\mathcal{P}$ adopts a symmetric encoder–decoder architecture composed of residual blocks and multi-scale skip connections, which facilitate effective information propagation across different spatial resolutions.
+All input images are resized to $256\times 256$, and the network progressively aggregates local and global contextual features through hierarchical downsampling and upsampling operations.
+To enable adaptive denoising under varying noise levels, $\mathcal{P}$ incorporates a learnable sigma embedding that conditions intermediate feature transformations, allowing the model to dynamically adjust its denoising behavior.
+We optimize $\mathcal{P}$ using the Adam optimizer with a learning rate of $1\times 10^{-4}$, a batch size of 8, and a total of 8 training epochs.
+Notably, $\mathcal{P}$ remains computationally efficient, containing only approximately 10 MB of parameters.
+A.2.2 Details of training residual mapper
+Figure A.3: Structure of the residual smoothness mapper.
+The structure of the proposed residual smoothness mapper $\mathcal{M}$ is illustrated in Figure A.3.
 Each block is designed to refine the purified feature representation while preserving its statistical structure and ensuring that the refinement magnitude adapts smoothly to the injected noise magnitude $\sigma$.
@@
 LayerNorm and FiLM($\sigma$). Each block begins by normalizing the input feature and injecting noise-level conditioning:
-$\tilde{{\bm{z}}}_{i}=\mathrm{LN}(\tilde{{\bm{z}}}_{i}),\hskip 28.80008pt[\gamma(\sigma),\beta(\sigma)]=\mathrm{FiLM}(\sigma),$ (S.28)
+$\tilde{{\bm{z}}}_{i}=\mathrm{LN}(\tilde{{\bm{z}}}_{i}),\qquad[\gamma(\sigma),\beta(\sigma)]=\mathrm{FiLM}(\sigma),$ (A.28)
 where the FiLM module takes the scalar noise strength $\sigma$ and outputs channel-wise affine parameters through a lightweight MLP:
-$\mathrm{FiLM}(\sigma)=W_{2}\,\phi(W_{1}\sigma)\in\mathbb{R}^{2D},$ (S.29)
+$\mathrm{FiLM}(\sigma)=W_{2}\,\phi(W_{1}\sigma)\in\mathbb{R}^{2D},$ (A.29)
 with $\phi$ denoting a GELU activation.
@@
 The outputs are fused as:
-$h_{i}=h_{\text{attn}}+h_{\text{mlp}}+0.5\,h_{\text{conv}},$ (S.30)
+$h_{i}=h_{\text{attn}}+h_{\text{mlp}}+0.5\,h_{\text{conv}},$ (A.30)
 where the conv branch is down-weighted following design practices for local/global feature fusion.
@@
 $\alpha_{i}=\mathrm{softplus}\!\left(\mathrm{MLP}_{\alpha}([\sigma,\log\sigma])\right),$
-$\sigma^{\beta_{i}}=\sigma^{\mathrm{softplus}(\mathrm{MLP}_{\beta}(\sigma))}.$ (S.31)
+$\sigma^{\beta_{i}}=\sigma^{\mathrm{softplus}(\mathrm{MLP}_{\beta}(\sigma))}.$ (A.31)
 Here, $\alpha_{i}$ controls the amplitude of the refinement, while $\beta_{i}$ modulates the exponent of the noise term, enabling nonlinear noise–feature interactions.
 The residual update is then computed as
-$\Delta{\bm{z}}_{i}=\left[(1+\gamma(\sigma))\,\bigl(v_{i}\cdot\alpha_{i}\sigma^{\beta_{i}}\bigr)+\beta(\sigma)\right]\odot\mathrm{Scale}_{i},$ (S.32)
+$\Delta{\bm{z}}_{i}=\left[(1+\gamma(\sigma))\,\bigl(v_{i}\cdot\alpha_{i}\sigma^{\beta_{i}}\bigr)+\beta(\sigma)\right]\odot\mathrm{Scale}_{i},$ (A.32)
 where $\gamma(\sigma)$ and $\beta(\sigma)$ are FiLM-generated per-channel affine parameters, and $\mathrm{Scale}_{i}$ is a learnable per-channel damping factor initialized to $5\times 10^{-4}$ for stability.
 The next feature is then updated via
-$\tilde{{\bm{z}}}_{i+1}=\tilde{{\bm{z}}}_{i}+\Delta{\bm{z}}_{i}.$ (S.33)
+$\tilde{{\bm{z}}}_{i+1}=\tilde{{\bm{z}}}_{i}+\Delta{\bm{z}}_{i}.$ (A.33)
 Overall effect of the mapper. The residual smoothness mapper thus provides the following benefits:
@@
 Collectively, $\mathcal{M}$ significantly enhances the smoothness and robustness of purified representations while avoiding distributional drift.
-Table S.6: Experimental results on adversarial robustness of different defense methods and MLLMs on image classification tasks.
-The Mapper $\mathcal{M}$ is a lightweight U-Net.
-The overall best results are shown in bold, and the best results without smoothing are underlined.
-Model Method M-Attack [23] FOA [14] AttackVLM [51]
-[…]
-LLaVA- 1.5-7B Org. 0.427 8.2% 78.2% 0.437 3.8 % 81.0 % 0.458 6.0% 78.4%
-Smoothed org. 0.590 84.8% 0.2% 0.600 87.2% 0.2% 0.605 87.0% 0.2%
-FARE [37] 0.574 55.4% 14.6% 0.521 24.8% 39.8% 0.508 25.8% 45.0%
-Smoothed FARE 0.695 75.6% 5.2% 0.681 66.2% 1% 0.671 68.0% 1%
-TeCoA [29] 0.731 67.4% 2% 0.626 29.0% 20.0% 0.592 33.6% 17.8%
-Smoothed TeCoA 0.790 70.6% 0.6% 0.714 58.8% 0.8% 0.712 56.4% 0.4%
-CLIP -L14 Org. 0.427 18.0% 37.2% 0.437 10.0 % 52.4 % 0.458 16.0% 48.4%
-Smoothed org. 0.590 88.0% 0.4% 0.600 90.0% 0.4% 0.604 90.4% 0.4%
-FARE [37] 0.574 70.4% 5.2% 0.521 42.0% 14.4% 0.508 35.6% 24.4%
-Smoothed FARE 0.695 84.4% 0% 0.682 78.4% 0% 0.671 76.4% 0%
-TeCoA [29] 0.731 81.2% 2.4% 0.626 39.2% 20.2% 0.592 35.8% 24.4%
-Smoothed TeCoA 0.790 87.6% 1.2% 0.714 75.6% 2.4% 0.712 73.2% 2.0%
-Figure S.4: Illustration of the evaluation process.
 Training details.
@@
 The training process employs the AdamW optimizer with an initial learning rate of $2\times 10^{-4}$ and a cosine annealing learning rate schedule.
-S.9 More experimental results
-S.9.1 Implementation details for main results
+Appendix A.3 More Details on Prediction-Wise Certification
+A.3.1 Preliminaries and Notation
+Let $f_{e}:\mathcal{X}\to\mathbb{R}^{d}$ be a (deterministic) feature encoder and let $\hat{f}_{e}$ denote its smoothed version produced by Feature-space Smoothing (FS).
+Given a clean input ${\bm{x}}\in\mathcal{X}$ and an adversarial input ${\bm{x}}^{\prime}\in\mathcal{X}$ satisfying $\|{\bm{x}}^{\prime}-{\bm{x}}\|\leq\epsilon$, define the clean feature and the adversarial (smoothed) feature as
+${\bm{z}}:=f_{e}({\bm{x}})\in\mathbb{R}^{d},\qquad\hat{\bm{z}}^{\prime}:=\hat{f}_{e}({\bm{x}}^{\prime})\in\mathbb{R}^{d}.$ (A.34)
+We further define their $\ell_{2}$-normalized counterparts:
+${\bm{u}}:=\frac{{\bm{z}}}{\|{\bm{z}}\|_{2}}\in\mathbb{S}^{d-1},\qquad{\bm{v}}:=\frac{\hat{\bm{z}}^{\prime}}{\|\hat{\bm{z}}^{\prime}\|_{2}}\in\mathbb{S}^{d-1},$ (A.35)
+where $\mathbb{S}^{d-1}:=\{{\bm{t}}\in\mathbb{R}^{d}:\|{\bm{t}}\|_{2}=1\}$ is the unit sphere. Thus, $\mathrm{Cos}(\hat{f}_{e}({\bm{x}}^{\prime}),f_{e}({\bm{x}}))={\bm{u}}^{\top}{\bm{v}}$.
+FS-certified feature cosine similarity.
+From Theorem 3.2, for any $\bm{x}$$^{\prime}$ with $\|{\bm{x}}^{\prime}-{\bm{x}}\|\leq\epsilon$, the cosine similarity between $\hat{f}_{e}({\bm{x}}^{\prime})$ and $f_{e}({\bm{x}})$ is lower bounded as
+${\bm{u}}^{\top}{\bm{v}}\;=\;\mathrm{Cos}\!\big(\hat{f}_{e}({\bm{x}}^{\prime}),f_{e}({\bm{x}})\big)\;\geq\;\gamma({\bm{x}};\epsilon),$ (A.36)
+where we denote the Feature Cosine Similarity Bound (FCSB) by
+$\gamma({\bm{x}};\epsilon)\;:=\;2\Phi\!\left(\Phi^{-1}(\hat{S}({\bm{x}}))-\epsilon\right)-1\in[-1,1].$ (A.37)
+For brevity, we write $\gamma:=\gamma({\bm{x}};\epsilon)$.
+Cosine-prototype classification head.
+We consider a cosine-based prototype classifier on the normalized feature ${\bm{t}}\in\mathbb{S}^{d-1}$.
+Let $\{{\bm{p}}_{k}\}_{k=1}^{K}\subset\mathbb{R}^{d}$ be learnable class prototypes satisfying $\|{\bm{p}}_{k}\|_{2}=1$ for all $k$.
+The class score (logit) for class $k$ is defined as
+$s_{k}({\bm{u}}):={\bm{u}}^{\top}{\bm{p}}_{k},$ (A.38)
+and the predicted label is
+$\hat{y}({\bm{u}}):=\arg\max_{k\in\{1,\dots,K\}}s_{k}({\bm{u}})\;=\;\arg\max_{k}{\bm{u}}^{\top}{\bm{p}}_{k}.$ (A.39)
+Let $y:=\hat{y}({\bm{u}})$ denote the clean prediction under $\bm{u}$}.
+A.3.2 Tight Cosine Similarity Bound
+Lemma A.3.1 (Tight bounds under a cosine constraint).
+Let ${\bm{u}},{\bm{v}},{\bm{p}}\in\mathbb{S}^{d-1}$ be unit vectors and suppose that ${\bm{u}}^{\top}{\bm{v}}\geq\gamma$ for some $\gamma\in[-1,1]$.
+Define $a:={\bm{u}}^{\top}{\bm{p}}\in[-1,1]$.
+Then the inner product ${\bm{v}}^{\top}{\bm{p}}$ is bounded as
+$\gamma a-\sqrt{1-\gamma^{2}}\sqrt{1-a^{2}}\;\leq\;{\bm{v}}^{\top}{\bm{p}}\;\leq\;\gamma a+\sqrt{1-\gamma^{2}}\sqrt{1-a^{2}}.$ (A.40)
+Proof sketch.
+The stated bounds are a standard consequence of spherical geometry.
+Since ${\bm{u}},{\bm{v}},{\bm{p}}\in\mathbb{S}^{d-1}$, we may interpret their inner products as cosines of the corresponding pairwise angles.
+In particular, let
+$a={\bm{u}}^{\top}{\bm{p}}=\cos\alpha,\qquad\gamma={\bm{u}}^{\top}{\bm{v}}=\cos\Delta,$
+where $\alpha,\Delta\in[0,\pi]$ denote the angles between $({\bm{u}},{\bm{p}})$ and $({\bm{u}},{\bm{v}})$, respectively.
+Under the constraint ${\bm{u}}^{\top}{\bm{v}}\geq\gamma$, the adversarial feature direction $\bm{v}$} is restricted to lie within a spherical cap centered at $\bm{u}$} with angular radius $\Delta$.
+Consequently, the angle between $\bm{v}$} and $\bm{p}$} lies in the interval $\alpha-\Delta,\;\alpha+\Delta]$.
+It follows that
+${\bm{v}}^{\top}{\bm{p}}=\cos(\alpha\pm\Delta)=\cos\alpha\cos\Delta\mp\sin\alpha\sin\Delta,$
+which yields the bounds in equation A.40 by substituting $\cos\alpha=a$ and $\cos\Delta=\gamma$.
+The bounds are tight and are attained when the three vectors ${\bm{u}},{\bm{v}},{\bm{p}}$ lie in a common two-dimensional subspace and $\bm{v}$} is rotated toward or away from $\bm{p}$} along the geodesic on the unit sphere.
+∎
+A.3.3 Closed-form Propagation to Prediction-level Certification
+We now use Lemma A.3.1 to certify that the predicted label remains unchanged under any adversarial perturbation within radius $\epsilon$.
+For the clean normalized feature $\bm{u}$}, define the clean class scores
+$a_{k}:=s_{k}({\bm{u}})={\bm{u}}^{\top}{\bm{p}}_{k},\qquad k=1,\dots,K.$ (A.41)
+Let $y:=\arg\max_{k}a_{k}$ be the clean predicted class.
+Theorem A.3.2 (Prediction-level certification via FS with cosine prototypes).
+Fix a clean input $\bm{x}$} and let $\gamma=\gamma({\bm{x}};\epsilon)$ be the FS-derived FCSB from Theorem 3.2.
+For any adversarial input $\bm{x}$$^{\prime}$ satisfying $\|{\bm{x}}^{\prime}-{\bm{x}}\|\leq\epsilon$ and each class $k$, we have:
+$s_{k}({\bm{v}})={\bm{v}}^{\top}{\bm{p}}_{k}$ $\leq\gamma a_{k}+\sqrt{1-\gamma^{2}}\sqrt{1-a_{k}^{2}},$ (A.42)
+$s_{k}({\bm{v}})={\bm{v}}^{\top}{\bm{p}}_{k}$ $\geq\gamma a_{k}-\sqrt{1-\gamma^{2}}\sqrt{1-a_{k}^{2}}.$ (A.43)
+In particular, if the following certified margin condition holds:
+$\underbrace{\gamma a_{y}-\sqrt{1-\gamma^{2}}\sqrt{1-a_{y}^{2}}}_{=:~\mathrm{LB}_{y}(\gamma)}\;>\;\max_{k\neq y}\underbrace{\Big(\gamma a_{k}+\sqrt{1-\gamma^{2}}\sqrt{1-a_{k}^{2}}\Big)}_{=:~\mathrm{UB}_{k}(\gamma)},$ (A.44)
+then the classifier prediction is provably invariant within the $\epsilon$-ball:
+$\arg\max_{k}s_{k}({\bm{v}})=y\quad\text{for all }{\bm{x}}^{\prime}\text{ such that }\|{\bm{x}}^{\prime}-{\bm{x}}\|\leq\epsilon.$ (A.45)
+Proof.
+By Theorem 3.2, for any $\bm{x}$$^{\prime}$ with $\|{\bm{x}}^{\prime}-{\bm{x}}\|\leq\epsilon$, we have
+$\mathrm{Cos}\!\big(\hat{f}_{e}({\bm{x}}^{\prime}),f_{e}({\bm{x}})\big)={\bm{u}}^{\top}{\bm{v}}\geq\gamma.$ (A.46)
+Fix any class $k$.
+Since ${\bm{u}},{\bm{v}},{\bm{p}}_{k}\in\mathbb{S}^{d-1}$ and ${\bm{u}}^{\top}{\bm{v}}\geq\gamma$, applying Lemma A.3.1 with ${\bm{p}}={\bm{p}}_{k}$ and $a=a_{k}={\bm{u}}^{\top}{\bm{p}}_{k}$ yields the per-class upper and lower bounds equation A.42–equation A.43.
+Now consider the clean predicted class $y=\arg\max_{k}a_{k}$. From equation A.43, the adversarial score of class $y$ is bounded from below as
+$s_{y}({\bm{v}})\geq\mathrm{LB}_{y}(\gamma).$ (A.47)
+For any other class $k\neq y$, from equation A.42 we have
+$s_{k}({\bm{v}})\leq\mathrm{UB}_{k}(\gamma).$ (A.48)
+If the certified margin condition equation A.44 holds, then for all $k\neq y$,
+$s_{y}({\bm{v}})\;\geq\;\mathrm{LB}_{y}(\gamma)\;>\;\mathrm{UB}_{k}(\gamma)\;\geq\;s_{k}({\bm{v}}),$ (A.49)
+which implies $s_{y}({\bm{v}})>s_{k}({\bm{v}})$ for every $k\neq y$. Therefore, the argmax prediction under $\bm{v}$} remains $y$, i.e.,
+$\arg\max_{k}s_{k}({\bm{v}})=y.$ (A.50)
+This holds for any $\bm{x}$$^{\prime}$ with $\|{\bm{x}}^{\prime}-{\bm{x}}\|\leq\epsilon$, which concludes the proof. ∎
+Appendix A.4 More Experimental Results
+Figure A.4: Illustration of the evaluation process.
+A.4.1 Implementation details for main results
 Image captioning.
@@
 The adversarial objective is to induce the MLLM to incorrectly choose “None of the above”.
-An illustration of the implementation process is shown in Figure S.4.
-S.9.2 Experimental results of using a lightweight U-net as noise purifier
-The comparative results of different defense methods for image classification are summarized in Table S.6.
-In this experiment, the original diffusion-based noise purifier is replaced with a lightweight U-Net denoiser containing approximately 43.7M parameters.
-The results show that with this compact denoiser, FS-PSM consistently achieves strong and superior robustness, outperforming adversarial training–based baselines by a substantial margin.
-Furthermore, integrating FS-PSM with existing defenses yields additional robustness gains, leading to consistently favorable improvements across all evaluated settings.
-These results collectively demonstrate the effectiveness and generalizability of the proposed method.
-S.9.3 Experimental results on large attack bound
-Table S.7: Experimental results on adversarial robustness of different defense methods and MLLMs on image classification tasks. The attack bound $\epsilon=32/255$.
-Model Method FOA [14]
+An illustration of the implementation process is shown in Figure A.4.
+A.4.2 Experimental results on large attack bound
+Table A.8: Experimental results on adversarial robustness of different defense methods and MLLMs on image classification tasks. The attack bound $\epsilon=32/255$.
+Model Method FOA (Jia et al., 2025)
 […]
@@
 Smoothed org. 0.512 35.2% 24.2%
-FARE [37] 0.408 10.4% 64.2%
+FARE (Schlarmann et al., 2024) 0.408 10.4% 64.2%
 Smoothed FARE 0.605 48.4% 19.2%
-TeCoA [29] 0.454 12.0% 29.6%
+TeCoA (Mao et al., 2023) 0.454 12.0% 29.6%
 Smoothed TeCoA 0.576 32.0% 9.6%
-The experimental results under a large adversarial perturbation of $\varepsilon=32/255$ are presented in Table S.7.
-In addition, Figure S.5 provides visualizations of the adversarial examples generated by different attacks and perturbation bounds.
+The experimental results under a large adversarial perturbation of $\varepsilon=32/255$ are presented in Table A.8.
+In addition, Figure A.5 provides visualizations of the adversarial examples generated by different attacks and perturbation bounds.
 These results show that nearly all existing defenses show great performance drop under such a strong attack.
-Under this challenging setting, the best accuracy is obtained by integrating FS-PSM with FARE, reaching 48.4% on LLaVA, while the lowest ASR (9.6%) is achieved by combining FS-PSM with TeCoA.
+Under this challenging setting, the best accuracy is obtained by integrating FS-GSB with FARE, reaching 48.4% on LLaVA, while the lowest ASR (9.6%) is achieved by combining FS-GSB with TeCoA.
 These findings further highlight the effectiveness and adaptability of the proposed framework, even against large-magnitude adversarial perturbations.
-Figure S.5: Visualization of the adversarial examples.
-S.9.4 Analysis of robustness-efficiency trade-off
-Table S.8: Analysis of the robustness-efficiency trade-off of using different $n_{0}$.
-Model $n_{0}$ Robustness Efficiency
-[…]
-LLaVA- 1.5-7B (diffusion purifier) 1 0.595 86.2% 0.6% 0.28
-2 0.652 88.4% 0.4% 0.41
-4 0.697 88.8% 0.4% 0.66
-8 0.719 89.6% 0.4% 1.12
-16 0.728 89.2% 0.4% 2.10
-LLaVA- 1.5-7B (lightweight purifier) 1 0.494 84.4% 0.2% 0.20
-2 0.565 85.4% 0% 0.24
-4 0.600 87.2% 0.2% 0.32
-8 0.623 87.8% 0.4% 0.49
-16 0.643 86.8% 0.4% 0.84
-The robustness–efficiency trade-off results are summarized in Table S.8.
-In this experiment, we adopt the strong FOA attack as the adversary and report the average inference time per sample on LLaVA-1.5-7B using a single RTX 4090 GPU.
-The results show that increasing the sampling number $n_{0}$ consistently improves the FCS under adversarial perturbations.
-Specifically, increasing $n_{0}$ from 0 to 16 raises FCS from approximately 0.595 to 0.728, which aligns with our theoretical analysis in Section 3.
-However, this robustness gain comes at a substantial computational cost.
-For example, increasing $n_{0}$ from 4 to 16 raises the inference latency from 0.66 seconds to 2.10 seconds per sample, revealing an inherent trade-off between robustness and efficiency.
+Figure A.5: Visualization of the adversarial examples.